### v2

Abstract
Video-to-video editing involves editing a source video along with additional control (such as text prompts, subjects, or styles) to generate a new video that aligns with the source video and the provided control. Traditional methods have been constrained to certain editing types, limiting their ability to meet the wide range of user demands. In this paper, we introduce AnyV2V, a novel training-free framework designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model (e.g. InstructPix2Pix, InstantID, etc) to modify the first frame, (2) utilizing an existing image-to-video generation model (e.g. I2VGen-XL) for DDIM inversion and feature injection. In the first stage, AnyV2V can plug in any existing image editing tools to support an extensive array of video editing tasks. Beyond the traditional prompt-based editing methods [20, 55, 15, 23, 36, 54], AnyV2V also can support novel video editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods. In the second stage, AnyV2V can plug in any existing image-to-video models to perform DDIM inversion and intermediate feature injection to maintain the appearance and motion consistency with the source video. On the prompt-based editing, we show that AnyV2V can outperform the previous best approach by 35% on prompt alignment, and 25% on human preference. On the three novel tasks, we show that AnyV2V also achieves a high success rate. We believe AnyV2V will continue to thrive due to its ability to seamlessly integrate the fast-evolving image editing methods. Such compatibility can help AnyV2V to increase its versatility to cater to diverse user demands.
Keywords:
video editing diffusion models plug-and-play
Figure 1: We build a universal framework AnyV2V to handle all types of video-to-video editing tasks conditioned on different inputs. AnyV2V disentangles the video editing process into two stages: (1) first-frame image editing and (2) image-to-video generation with DDIM inversion + feature injection.
1 Introduction
The task of video-to-video editing represents a crucial area of research that enables users to edit and create videos according to their preferences seamlessly. In this domain, an AI model is tasked with processing a source video along with various forms of guidance—such as text prompts, facial identity, styles, subjects, depth, and more—to synthesize a new video. It is imperative that the synthesized video not only remains faithful to the source video but also accurately incorporates the additional guidance provided. This requirement ensures that the synthesized video meets the specific desires and intentions of the user. In order to tackle the video-to-video generation tasks, an extensive range of methods [20, 15, 54, 55, 3, 36, 35, 23, 15] have been proposed. These methods are built on top of the open-source image or video diffusion models [45, 17, 62, 9, 53]. However, these methods are all limited to performing specific types of editing tasks. For instance, TokenFlow [20] excels at text-guided style transfer but falls short in localized editing. VideoP2P [36] are restricted to follow certain types of editing prompts. Besides, some of these approaches [55, 23, 17] require additional parameter tuning or video feature extraction (e.g. tracking, point correspondence, etc) to ensure appearance or temporal consistency. Such overhead could lead to much higher time and memory costs. These limitations make them incapable of meeting the diverse user requirements, necessitating the establishment of separate video editing workflows for different tasks.
In recognition of these limitations, a compelling need arises to conceptualize and develop a simple and general-purpose video editing solution. In this paper, we introduce a unified yet simple framework, AnyV2V, designed to address (possibly) any video editing tasks. Our primary insight is that any video editing task can be effectively decomposed into two pivotal stages:
  1. 1.
    Apply an off-the-shelf task-specific image editing model over the first frame.
  2. 2.
    Apply an image-to-video model to perform DDIM inversion on the source video and then inject the intermediate features to generate the new video.
In the first stage, AnyV2V employs an existing task-specific image editing method to edit the first frame(Section 4.1). In the second stage, we aim to propagate the edited effect across the entire video while ensuring alignment with the source video. To achieve this, we follow an invert-and-generate framework [56, 38]. In contrast to prior methods that employed text-to-video models [3, 4] for DDIM inversion [47], we employ the image-to-video models [62, 12, 44] for DDIM inversion(Section 4.2) to enable the first-frame condition. With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model is able to generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video. To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in convolution layers, spatial attention layers and temporal attention layers in the denoising U-Net of the I2V model. These techniques are discussed in detail in Section 4.3 and Section 4.4, and we demonstrate their effectiveness by conducting an ablation study in Section 5.5. By performing the two-stage editing process, AnyV2V effectively offloads the editing operation to existing image editing tools. This disentanglement helps AnyV2V excel in:
  1. 1.
    Compatibility: AnyV2V has superior compatibility with all the image editing methods. We demonstrate that AnyV2V can seamlessly build on top of advanced image editing methods, such as InstructPix2Pix [6], InstantID [51], NST [18], AnyDoor [11] to perform diverse types of editing.
  2. 2.
    Simplicity: AnyV2V is a tuning-free approach without requiring any additional video features to achieve high appearance and temporal consistency.
In this paper, we showcase AnyV2V’s versatility and effectiveness by the qualitative and quantitative results. We comprehensively evaluate our method on four important tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing and (4) identity manipulation. We collected a new dataset of roughly 100 examples based on the four tasks to evaluate the performance of AnyV2V and other baselines [20, 15, 55]. We show that AnyV2V is among the first to perform (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation in the video domain. Moreover, the AnyV2V exhibits 25% human preference improvement over the previous best model [20] in (1) prompt-based editing. We also perform a comprehensive ablation study to show the impact of our design choices. Our contributions are summarized as follows:
  1. 1.
    We introduce AnyV2V, a plug-and-play unified framework tailored for a diverse range of video-to-video editing tasks.
  2. 2.
    To the best of our knowledge, we are the first to perform video editing using pre-trained I2V models, marking a novel paradigm in this field.
  3. 3.
    Universal compatibility with image editing methods. AnyV2V is compatible with any image editing method, extending any image editing method into the video domain at no cost.
  4. 4.
    We show, both quantitatively and qualitatively, that our method outperforms existing state-of-the-art baselines on prompt-based editing and exhibits robust performance on three novel tasks: reference-based style transfer, subject-driven editing, and identity manipulation.
2 Related Work
2.0.1 Text-to-Video (T2V) Generation.
Recent advancements in diffusion models [27, 46] have led to a huge development in text-to-video (T2V) generation. Existing diffusion-based T2V models generally follow the paradigm of extending text-to-image (T2I) generation models to the video domain, where additional temporal modelling layers are added to the T2I model to enable video generation. Recent T2V generation models are developed from pixel-space diffusion models [29, 26, 19], latent diffusion models (LDMs) [63, 1, 5, 24, 8, 53], as well as diffusion transformers [25, 10, 41]. LDM-based [45] T2V generation models are the most extensively studied methods due to their high efficiency and the open-sourced Stable Diffusion [45] models. T2V models have been shown to be foundational models for a wide range of applications such as I2V generation [57], personalized video generation [39] and video editing [3].
2.0.2 Image-to-Video (I2V) Generation.
To achieve more precise control over the video generation process, a series of I2V generation methods employing additional image conditions in the video generation process have emerged. Earlier efforts such as VideoComposer [52] and VideoCrafter1 [8] use the input image as a style reference and cannot fully preserve the visual details in the output videos. More recently, methods such as DynamiCrafter [57], I2VGen-XL [62], SEINE [12] and ConsistI2V [44] have overcome this problem and can generate videos that are more consistent with the input image (often used as the first frame of the video). I2V generation models have also been employed in a wide range of applications such as long video generation [12, 59, 44], generative frame interpolation [57, 12] and video storytelling [57, 59]. However, to the best of our knowledge, no methods have applied I2V generation models for video editing. We aim to investigate this area and propose a highly flexible training-free framework for video editing using I2V generation models.
2.0.3 Semantic Image and Video Editing.
Visual generation has attracted considerable attention within the field; however, visual manipulation also represents a significant and popular area of interest. The research area of image editing encompasses a wide range of tasks. The popular tasks include stylization [18, 13, 37], identity manipulation [51], subject-driven editing [34, 22, 11], and localized editing [40, 2, 48, 33]. One interesting direction is to train a model to understand text inputs to perform general image editing tasks [60, 6, 49]. However, image editing remained an open and challenging problem as none of the work was able to tackle all the tasks effectively [31]. On the other hand, the video editing problem was often perceived as an extension of image editing problems, incorporating the time dimension. Recent video editing works [20, 54, 42, 15, 58, 7] employ text prompts to execute video editing tasks. These approaches have opened new avenues for interactive and intuitive video manipulation, allowing users to guide the editing process through natural language. However, such methodologies fail to offer precise control to the user, often resulting in edits that may not fully align with the user’s intentions or the desired level of detail due to the ambiguity of natural language and the limit of model capability. Our work is the first to enable precise video editing using pre-trained Image-to-Video (I2V) models. More importantly, our paradigm merges the two problems into a single domain to create a unified solution, offering high controllability in video editing tasks.
3 Preliminary
3.0.1 Image-to-Video (I2V) Generation Models.
In this work, we focus on leveraging latent diffusion-based [45] I2V generation models for video editing. Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$. The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (1)
$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V,$ (2)
where $z$ is the input hidden state to the self-attention layer and $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable projection matrices that map $z$ onto query, key and value vectors, respectively. For spatial self-attention, $z$ represents a sequence of spatial tokens from each frame. For temporal self-attention, $z$ is composed of tokens located at the same spatial position across all frames.
3.0.2 DDIM Inversion.
The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM [47] sampling algorithm. The reverse process of DDIM sampling, known as DDIM inversion [38, 16], allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
3.0.3 Plug-and-Play (PnP) Diffusion Features.
Tumanyan et al. [49] proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g. legs or torso of a human body) during the image generation process.
Given an input source image $I^{S}$ and a target prompt $P$, PnP first performs DDIM inversion to obtain the image’s corresponding noise $\{\mathbf{z}^{S}_{t}\}_{t=1}^{T}$ at each time step $t$. It then collects the convolution features $\{f^{l}_{t}\}$ and attention scores $\{A^{l}_{t}\}$ from some predefined layers $l$ at each time step $t$ of the backward diffusion process $\mathbf{z}_{t-1}^{S}=\epsilon_{\theta}(\mathbf{z}_{t}^{S},\varnothing,t)$, where $\varnothing$ denotes the null text prompt during denoising.
To generate the edited image $I^{*}$, PnP starts from the initial noise of the source image (i.e. $\mathbf{z}_{T}^{*}=\mathbf{z}_{T}^{S}$) and performs feature injection during denoising: $\mathbf{z}_{t-1}^{*}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},P,t,\{f^{l}_{t},A^{l}_{t}\})$, where $\epsilon_{\theta}(\cdot,\cdot,\cdot,\{f^{l}_{t},A^{l}_{t}\})$ represents the operation of replacing the intermediate feature and attention scores $\{f^{l*}_{t},A^{l*}_{t}\}$ with $\{f^{l}_{t},A^{l}_{t}\}$. This feature injection mechanism ensures $I^{*}$ to preserve the layout and structure from $I^{S}$ while reflecting the description in $P$. To control the feature injection strength, PnP also employs two thresholds $\tau_{f}$ and $\tau_{A}$ such that the feature and attention scores are only injected in the first $\tau_{f}$ and $\tau_{A}$ denoising steps. Our method extends this feature injection mechanism to I2V generation models, where we inject features in convolution, spatial, and temporal attention layers. We show the detailed design of AnyV2V in Section 4.
4 AnyV2V
Figure 2: AnyV2V framework. Our framework takes a source video $V^{S}$ as input. In the first stage, we apply a block-box image editing method on the first frame $I_{1}$ according to the editing task. In the second stage, the source video is inverted to initial noise $z_{T}^{S}$, which is then denoised using DDIM sampling. During the sampling process, we extract spatial features, spatial attention, and temporal attention from the image-to-video models’ decoder layers. To generate our edited video, we perform a DDIM sampling by fixing $z_{T}^{*}$ as $z_{T}^{T}$ and use the edited first frame as the conditional signal. During the sampling, we inject the features and attention into corresponding layers of the model.
Our method presents a two-stage approach to video editing. Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ dentes the video length. We extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$, where $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc. In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$ and the semantic information represented in the edited first frame $I^{*}_{1}$ and the target prompt $\mathbf{s^{*}}$. An overall illustration of our video editing pipeline is shown in Figure 2. In this section, we explain each core component of our method.
4.1 Flexible First Frame Editing
In visual manipulation, controllability is a key element in performing precise editing. Current video editing models lack this controllability as most of the existing methods can only edit videos based on target text prompts (e.g. TokenFlow [20]) or text instructions (e.g. InsV2V [14]). Our AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame. This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [18, 21, 37], mask-based image editing [40, 2], image inpainting [48, 33], identity-preserving image editing [51], and subject-driven image editing [11, 34, 22]. This stage is highly flexible and it can even be done by human experts.
4.2 Structural Guidance using DDIM Inverison
To ensure the generated videos from the I2V generation model follow the general structure as presented in the source video, we employ DDIM inversion to obtain the latent noise of the source video at each time step $t$. Specifically, we perform the inversion without text prompt condition but with the first frame condition. Formally, given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, we obtain the inverted latent noise for time step $t$ as:
$\mathbf{z}^{S}_{t}=\mathrm{DDIM\_Inv}(\epsilon_{\theta}(\mathbf{z}_{t+1},I_{1},\varnothing,t)),$ (3)
[…]
4.3 Appearance Guidance via Spatial Feature Injection
Our empirical observation (cf. Section 5.5) suggests that I2V generation models already have some editing capabilities by only using the edited first frame and DDIM inverted noise as the model input. However, we find that this simple approach is often unable to correctly preserve the background in the edited first frame and the motion in the source video, as the conditional signal from the source video encoded in the inverted noise is limited.
To enforce consistency with the source video, we perform feature injection in both convolution layers and spatial attention layers in the denoising U-Net. During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$. We preserve two types of hidden features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers. We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors. We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance [28]. We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
4.4 Motion Guidance through Temporal Feature Injection
The spatial feature injection mechanism described in Section 4.3 significantly enhances the background and overall structure consistency of the edited video. While it also helps maintain the source video motion to some degree, we observe that the edited videos will still have a high chance of containing incorrect motion compared to the source video. On the other hand, we notice that I2V generation models, or video diffusion models in general, are often initialized from pre-trained T2I models and continue to be trained on video data. During the training process, parameters in the spatial layers are often frozen or set to a lower learning rate such that the pre-trained weights from the T2I model are less affected, and the parameters in the temporal layers are more extensively updated during training. Therefore, it is likely that a large portion of the motion information is encoded in the temporal layers of the I2V generation models. Concurrent work [3] also observes that features in the temporal layers show similar characteristics with optical flow [30], a pattern that is often used to describe the motion of the video.
To better reconstruct the source video motion in the edited video, we propose to also inject the temporal attention features in the video generation process. Similar to spatial attention injection, we collect the source video temporal self-attention queries $Q^{l_{3}}_{t}$ and keys $K^{l_{3}}_{t}$ from some U-Net decoder layers represented by $l_{3}$ and inject them into the edited video denoising branches. We also only apply temporal attention injection in the first $\tau_{ta}$ steps during sampling. Overall, combining the spatial and temporal feature injection mechanisms, the denoising process of our AnyV2V can be represented by:
$\mathbf{z}^{*}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},I^{*},\mathbf{s}^{*},t,\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\}).$ (4)
Our proposed spatial and temporal feature injection scheme enables tuning-free adaptation of I2V generation models for video editing. Our experimental results demonstrate that each component in our design is crucial to the accurate editing of source videos. We showcase more qualitative results for the effectiveness of our model components in Section 5.
5 Experiments
In this section, we first introduce four video editing tasks that we focus on evaluating our AnyV2V framework: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation. We then assess the effectiveness of AnyV2V both quantitatively and qualitatively on these four tasks. For (1) prompt-based editing, we compare AnyV2V with SoTA text-guided video editing baselines. For the three novel tasks (2)(3)(4), as these tasks require reference images to guide video editing, current text-based video editing methods are incapable of performing such edits. We thus compare the performance of AnyV2V using different I2V generation backbones.
5.1 Tasks Definition
  • 1.
    Prompt-based Editing: allows users to manipulate video content using only natural language. This can include descriptive prompts or instructions. With the prompt, Users can perform a wide range of edits, such as incorporating accessories, spawning or swapping objects, adding effects, or altering the background.
  • 2.
    Reference-based Style Transfer: In the realm of style transfer tasks, the artistic styles of Monet and Van Gogh are frequently explored, but in real-life examples, users might want to use a distinct style based on one particular artwork. In reference-based style transfer, we focus on using a style image as a reference to perform video editing. The edited video should capture the distinct style of the referenced artwork.
  • 3.
    Subject-driven Editing: In subject-driven video editing, we aim at replacing an object in the video with a target subject based on a given subject image while maintaining the video motion and persevering the background.
  • 4.
    Identity Manipulation: Identity manipulation allows the user to manipulate video content by replacing a person with another person’s identity in the video based on an input image of the target person.
5.2 Implementation Details
We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl. [62], ConsistI2V [44] and SEINE [12]. For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps. We use the DDIM [47] sampler and set $T$ to the default values of the selected I2V models. Following PnP [50], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections. During sampling, we apply text classifier-free guidance (CFG) [28] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits. We refer readers to the supplementary materials for more discussions on our implementation details and hyperparameter settings.
To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [6], style transfer model Neural Style Transfer (NST) [18], subject-driven image editing model AnyDoor [11], and identity-driven image editing model InstantID [51]. We experiment with only the successfully edited frames, which is crucial for our method.
5.3 Quantitative Evaluations
5.3.1 Prompt-based Editing.
For (1) prompt-based editing, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting. We compare AnyV2V against three baseline models: Tune-A-Video [55], TokenFlow [20] and FLATTEN [15]. Human evaluation results in Table 1 demonstrate that our model achieves the best overall preference and prompt alignment among all methods, and AnyV2V (I2VGen-XL) is the most preferred method. We conjecture that the gain is coming from our compatibility with state-of-the-art image editing models.
We also employ automatic evaluation metrics on our edited video of the human evaluation datasets. Following previous works [7, 3], our automatic evaluation employs the CLIP [43] model to assess both text alignment and temporal consistency. For text alignment, we calculate the CLIP-Score, specifically by determining the average cosine similarity between the CLIP text embeddings derived from the editing prompt and the CLIP image embeddings across all frames. For temporal consistency, we evaluate the average cosine similarity between the CLIP image embeddings of every pair of consecutive frames. These two metrics are referred to as CLIP-Text and CLIP-Image, respectively. Our automatic evaluations in Table 1 demonstrate that our model is competitive in prompt-based editing compared to baseline methods.
5.3.2 Reference-based Style Transfer; Identity Manipulation and Subject-driven Editing
For novel tasks (2),(3) and (4), we evaluate the performance of three I2V generation models using human evaluations and show the results in Table 2. As these tasks require reference images instead of text prompts, we focus on evaluating the reference alignment and overall preference of the edited videos. According to the results, we observe that AnyV2V (I2VGen-XL) is the best model across all tasks, underscoring its robustness and versatility in handling diverse video editing tasks. AnyV2V (SEINE) and AnyV2V (ConsistI2V) show varied performance across tasks. AnyV2V (SEINE) performs good reference alignment in reference-based style transfer and identity manipulation, but falls short in subject-driven editing with lower scores. On the other hand, AnyV2V (ConsistI2V) shines in subject-driven editing, achieving second-best results in both reference alignment and overall preference. Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing [32], we also report the image editing success rate in Table 2 to clarify that our method relies on a good image frame edit.
[…]
Task Prompt-based Editing
Method Human Evaluation $\uparrow$ CLIP Scores $\uparrow$
Alignment Overall CLIP-Text CLIP-Image
Tune-A-Video [55] 15.2% 2.1% 0.2902 0.9704
[…]
AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
[…]
Task Reference-based Subject-driven Identity
Style Transfer Editing Manipulation
Image Editing Method NST [18] AnyDoor [11] InstantID [51]
Image Editing Success Rate $\approx$90$\approx$10$\approx$80%
Human Evaluation Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$
[…]
AnyV2V (I2VGen-XL) 100.0% 76.9% 93.9% 84.8% 90.1% 45.4%
5.4 Qualitative Results
5.4.1 Prompt-based Editing.
By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing. In our experiment, we used InstructPix2Pix [6] for the first frame edit. Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video. Comparing our work with the three baseline models TokenFlow [20], FLATTEN [15], and Tune-A-Video [55], the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt. The color tone and object shapes are also tilted. It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes. The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modelling.
Figure 3: AnyV2V is robust in a wide range of prompt-based editing tasks while maintaining the background. The generated results align the most with the text prompt and also maintain high motion consistency.
Figure 4: With different image editing models, AnyV2V can achieve a wide range of editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation.
5.4.2 Reference-based Style Transfer.
Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST [18] to obtain the edited frame. This level of controllability offers artists the unprecedented opportunity to use their art as a reference for video editing, opening new avenues for creative expression. As demonstrated in Figure 4, our method captures the distinctive style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately, while such an edit is often hard to perform using existing text-guided video editing methods.
5.4.3 Subject-driven Editing.
In our experiment, we employed a subject-driven image editing model AnyDoor [11] for the first frame editing. AnyDoor allows replacing any object in the target image with the subject from only one reference image. We observe from Figure 4 that AnyV2V produces highly motion-consistent videos when performing subject-driven object swapping. In the first example, AnyV2V successfully replaces the cat with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video. In the second example, the car is replaced by our desired car while maintaining the rotation angle in the edited video.
5.4.4 Identity Manipulation.
By integrating the identity-preserved image personalization model InstantID [51] with ControlNet [61], this approach enables the replacement of an individual’s identity to create an initial frame. Our AnyV2V framework then processes this initial frame to produce an edited video, swapping the person’s identity as showcased in Figure 4. To the best of our knowledge, our work is the first to provide such flexibility in the video editing models. Note that the InstantID with ControlNet method will alter the background due to its model property. It is possible to leverage other identity-preserved image personalization models and apply them to AnyV2V to preserve the background.
Table 3: Ablation study results for AnyV2V (I2VGen-XL). T. Injection and S. Injection correspond to temporal and spatial feature injection mechanisms, respectively.
Model CLIP-Image $\uparrow$
AnyV2V (I2VGen-XL) 0.9648
AnyV2V (I2VGen-XL) w/o T. Injection 0.9652
AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection 0.9637
AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection & DDIM Inversion 0.9607
5.5 Ablation Study
To verify the effectiveness of our design choices, we conduct an ablation study by iteratively disabling the three core components in our model: temporal feature injection, spatial feature injection, and DDIM inverted latent as initial noise. We use AnyV2V (I2VGen-XL) and a subset of 20 samples in this ablation study and report both the frame-wise consistency results using CLIP-Image score in Table 3 and qualitative comparisons in Figure 5. We provide more ablation analysis of other design considerations of our model in the supplementary materials.
Figure 5: Visual comparisons of AnyV2V’s editing results after disabling temporal feature injection (T.I.), spatial feature injection (S.I.) and DDIM inverted initial noise (D.I.).
5.5.1 Effectiveness of Temporal Feature Injection.
[…]
5.5.2 Effectiveness of Spatial Feature Injection.
As shown in Table 3, we observe a drop in the CLIP-Image score after removing the spatial feature injection mechanisms from our model, indicating that the edited videos are not smoothly progressed across consecutive frames and contain more appearance and motion inconsistencies. Further illustrated in the third row of Figure 5, removing spatial feature injection will often result in incorrect subject appearance and pose (as shown in the “ballet dancing” case) and degenerated background appearance (evident in the “couple sitting” case). These observations demonstrate that directly generating edited videos from the DDIM inverted noise is often not enough to fully preserve the source video structures, and the spatial feature injection mechanisms are crucial for achieving better editing results.
5.5.3 DDIM Inverted Noise as Structural Guidance.
Finally, we observe a further decrease in CLIP-Image scores and a significantly degraded visual appearance in both examples in Figure 5 after replacing the initial DDIM inverted noise with random noise during sampling. This indicates that the I2V generation models become less capable of animating the input image when the editing prompt is completely out-of-domain and highlights the importance of the DDIM inverted noise as the structural guidance of the edited videos.
6 Conclusion
In this paper, we present AnyV2V, a new unified framework for video editing. Our framework is training-free, highly cost-effective, and can be applied to any I2V model. To perform video editing with high precision, we propose a two-stage approach to first edit the initial frame of the source video and then condition an image-to-video (I2V) model with the initial frame. Comprehensive experiments have shown that our method achieves outstanding outcomes across a broad spectrum of applications and shows a high level of controllability.
7 Limitations
7.0.1 Inaccurate Edit from Image Editing Models.
As our method relies on an initial frame edit, the image editing models are used. However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [32]. For example, in the subject-driven video editing task, we found that AnyDoor [11] requires several tries to get a good editing result. Efforts are required in manually picking a good edited frame. We expect that in the future better image editing models will minimize such effort.
7.0.2 Limited ability of I2V models.
We found that the results from our method cannot follow the source video motion if the motion is fast (e.g. billiard balls hitting each other at full speed) or complex (e.g. a person clipping her hair). One possible reason is that the current popular I2V models are generally trained on slow-motion videos, such that lacking the ability to regenerate fast or complex motion even with motion guidance. We anticipate that the presence of a robust I2V model can address this issue.
8 Potential Negative Impacts
Misinformation spread and Privacy violations. As our technique allows for object manipulation, it can produce highly realistic yet completely fabricated videos of one individual or subject. There is a risk that harmful actors could exploit our system to generate counterfeit videos to disseminate false information. Moreover, the ability to create convincing counterfeit content featuring individuals without their permission undermines privacy protections, possibly leading to the illicit use of a person’s likeness for harmful purposes and damaging their reputation. These issues are similarly present in DeepFake technologies. To mitigate the risk of misuse, one proposed solution is the adoption of unseen watermarking, a method commonly used to tackle such concerns in image generation.
Supplementary Material
Appendix 0.A Discussion on Model Implementation Details
When adapting our AnyV2V to various I2V generation models, we identify two sets of hyperparameters that are crucial to the final video editing results. They are (1) selection of U-Net decoder layers ($l_{1}$, $l_{2}$ and $l_{3}$) to perform convolution, spatial attention and temporal attention injection and (2) Injection thresholds $\tau_{conv}$, $\tau_{sa}$ and $\tau_{ta}$ that control feature injections to happen in specific diffusion steps. In this section, we provide more discussions and analysis on the selection of these hyperparameters.
0.A.1 U-Net Layers for Feature Injection
Figure 6: Visualizations of the convolution, spatial attention and temporal attention features during video sampling for I2V generation models’ decoder layers. We feed in the DDIM inverted noise to the I2V models such that the generated videos (first row) are reconstructions of the source video.
To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL [62], ConsistI2V [44] and SEINE [12]. Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e. average attention weights for all other tokens attending to the current token). The results are shown in Figure 6.
According to the figure, we observe that the intermediate convolution features from different I2V models show similar characteristics during video generation: earlier layers in the U-Net decoder produce features that represent the overall layout of the video frames and deeper layers capture the high-frequency details such as edges and textures. We choose to set $l_{1}=4$ for convolution feature injection to inject background and layout guidance to the edited video without introducing too many high-frequency details. For spatial and temporal attention scores, we observe that the spatial attention maps tend to represent the semantic regions in the video frames while the temporal attention maps highlight the foreground moving subjects (e.g. the running woman in Figure 6). One interesting observation for I2VGen-XL is that its spatial attention operations in deeper layers almost become hard attention, as the spatial tokens only attend to a single or very few tokens in each frame. We propose to inject features in decoder layers 4 to 11 ($l_{2}=l_{3}=\{4,5,...,11\}$) to preserve the semantic and motion information from the source video.
0.A.2 Ablation Analysis on Feature Injection Thresholds
We perform additional ablation analysis using different feature injection thresholds to study how these hyperparameters affect the edited video.
0.A.2.1 Effect of Spatial Injection Thresholds $\tau_{conv},\tau_{sa}$
We study the effect of disabling spatial feature injection or using different $\tau_{conv}$ and $\tau_{sa}$ values during video editing and show the qualitative results in Figure 7. We find that when spatial feature injection is disabled, the edited videos fail to fully adhere to the layout and motion from the source video. When spatial feature injection thresholds are too high, the edited videos are corrupted by the high-frequency details from the source video (e.g. textures from the woman’s down jacket in Figure 7). Setting $\tau_{conv}=\tau_{sa}=0.2T$ achieves a desired editing outcome for our experiments.
Source
$\tau_{c,s}=0$
$\tau_{c,s}=0.2T$
$\tau_{c,s}=0.5T$
$\tau_{c,s}=0.7T$
$\tau_{c,s}=T$
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/source/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/0/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/2/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/5/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/7/015
[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/10/015
$\tau_{c,s}=0$
$\tau_{c,s}=0.2T$
$\tau_{c,s}=0.5T$
$\tau_{c,s}=0.7T$
$\tau_{c,s}=T$
Figure 7: Hyperparameter study on spatial feature injection. We find that $\tau_{sa}=0.2T$ is the best setting for maintaining the layout and structure in the edited video while not introducing unnecessary visual details from the source video. $\tau_{c,s}$ represents $\tau_{conv}$ and $\tau_{sa}$. (Editing prompt: teddy bear running. Videos clickable in Adobe Acrobat.)
0.A.2.2 Effect of Temporal Injection Threshold $\tau_{ta}$
We study the hyperparameter of temporal feature injection threshold $\tau_{ta}$ in different settings and show the results in Figure 8. We observe that in circumstances where $\tau_{ta}<0.5T$ ($T$ is the total denoising steps), the motion guidance is too weak that it leads to only partly aligned motion with the source video, even though the motion itself is logical and smooth. At $\tau_{ta}>0.5T$, the generated video shows a stronger adherence to the motion but distortion occurs. We employ $\tau_{ta}=0.5T$ in our experiments and find that this value strikes the perfect balance on motion alignment, motion consistency, and video fidelity.
Source
$\tau_{ta}=0$
$\tau_{ta}=0.2T$
$\tau_{ta}=0.5T$
$\tau_{ta}=0.7T$
$\tau_{ta}=T$
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/src/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.2_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.5_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.7_stp50/frames/frame_015
[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp1.0_stp50/frames/frame_015
$\tau_{ta}=0$
$\tau_{ta}=0.2T$
$\tau_{ta}=0.5T$
$\tau_{ta}=0.7T$
$\tau_{ta}=T$
Figure 8: Hyperparameter study on temporal feature injection. We find that $\tau_{ta}=0.5T$ to be the optimal setting as it balances motion alignment, motion consistency, and video fidelity. (Editing prompt: darth vader walking. Videos clickable in Adobe Acrobat.)
0.A.3 Backbones Performance
Among the three backbones of our framework, We found that AnyV2V (I2VGen-XL) tends to be the most robust model both qualitatively and quantitatively. It has a good generalization ability to produce consistent motions in the video with high visual quality. AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its data used during the training stage, thus harming the visual quality. AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking. One interesting property of our framework is that although in some cases the motion does not fully align with the source video, it produces video with high motion consistency and reasonable motion that is highly similar to the source video, indicating the flexibility of how the video can be edited while maintaining the video fidelity.
Appendix 0.B Human Evaluation Dataset
Our human evaluation dataset contains in total of 89 samples. For prompt-based editing, we employed InstructPix2Pix [6] to compose the examples. Topics include swapping objects, adding objects, and removing objects. For subject-driven editing, we employed AnyDoor [11] to replace objects with reference subjects. For Neural Style Transfer, we employed NST [18] to compose the examples. For identity manipulation, we employed InstantID [51] to compose the examples. See Table 4 for the statistic.
Table 4: Number of entries for Video Editing Evaluation Dataset
Category Number of Entries Image Editing Models Used
Prompt-based Editing 45 InstructPix2Pix [6]
Reference-based Style Transfer 20 NST [18]
Identity Manipulation 13 InstantID [51]
Subject-driven Editing 11 AnyDoor [11]
Total 89

### v3

@@
-
+AnyV2V: A Tuning-Free Framework For Any Video-to-Video Editing Tasks
 Abstract
-Video-to-video editing involves editing a source video along with additional control (such as text prompts, subjects, or styles) to generate a new video that aligns with the source video and the provided control.
-Traditional methods have been constrained to certain editing types, limiting their ability to meet the wide range of user demands.
-In this paper, we introduce AnyV2V, a novel training-free framework designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model (e.g.
-InstructPix2Pix, InstantID, etc) to modify the first frame, (2) utilizing an existing image-to-video generation model (e.g.
-I2VGen-XL) for DDIM inversion and feature injection.
-In the first stage, AnyV2V can plug in any existing image editing tools to support an extensive array of video editing tasks.
-Beyond the traditional prompt-based editing methods [20, 55, 15, 23, 36, 54], AnyV2V also can support novel video editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods.
-In the second stage, AnyV2V can plug in any existing image-to-video models to perform DDIM inversion and intermediate feature injection to maintain the appearance and motion consistency with the source video.
-On the prompt-based editing, we show that AnyV2V can outperform the previous best approach by 35% on prompt alignment, and 25% on human preference.
-On the three novel tasks, we show that AnyV2V also achieves a high success rate.
-We believe AnyV2V will continue to thrive due to its ability to seamlessly integrate the fast-evolving image editing methods.
-Such compatibility can help AnyV2V to increase its versatility to cater to diverse user demands.
-Keywords:
-video editing diffusion models plug-and-play
-Figure 1: We build a universal framework AnyV2V to handle all types of video-to-video editing tasks conditioned on different inputs.
-AnyV2V disentangles the video editing process into two stages: (1) first-frame image editing and (2) image-to-video generation with DDIM inversion + feature injection.
+In the dynamic field of digital content creation using generative models, state-of-the-art video editing models still do not offer the level of quality and control that users desire.
+Previous works on video editing either extended from image-based generative models in a zero-shot manner or necessitated extensive fine-tuning, which can hinder the production of fluid video edits.
+Furthermore, these methods frequently rely on textual input as the editing guidance, leading to ambiguities and limiting the types of edits they can perform.
+Recognizing these challenges, we introduce AnyV2V, a novel tuning-free paradigm designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model to modify the first frame, (2) utilizing an existing image-to-video generation model to generate the edited video through temporal feature injection.
+AnyV2V can leverage any existing image editing tools to support an extensive array of video editing tasks, including prompt-based editing, reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods.
+AnyV2V can also support any video length.
+Our evaluation indicates that AnyV2V significantly outperforms other baseline methods in automatic and human evaluations by significant margin, maintaining visual consistency with the source video while achieving high-quality edits across all the editing tasks.
+Figure 1: AnyV2V is an evolving framework to handle all types of video-to-video editing tasks without any parameter tuning.
+AnyV2V disentangles video editing into two simpler problems: (1) Single image editing and (2) Image-to-video generation with video reference.
 1 Introduction
-The task of video-to-video editing represents a crucial area of research that enables users to edit and create videos according to their preferences seamlessly.
-In this domain, an AI model is tasked with processing a source video along with various forms of guidance—such as text prompts, facial identity, styles, subjects, depth, and more—to synthesize a new video.
-It is imperative that the synthesized video not only remains faithful to the source video but also accurately incorporates the additional guidance provided.
-This requirement ensures that the synthesized video meets the specific desires and intentions of the user.
-In order to tackle the video-to-video generation tasks, an extensive range of methods [20, 15, 54, 55, 3, 36, 35, 23, 15] have been proposed.
-These methods are built on top of the open-source image or video diffusion models [45, 17, 62, 9, 53].
-However, these methods are all limited to performing specific types of editing tasks.
-For instance, TokenFlow [20] excels at text-guided style transfer but falls short in localized editing.
-VideoP2P [36] are restricted to follow certain types of editing prompts.
-Besides, some of these approaches [55, 23, 17] require additional parameter tuning or video feature extraction (e.g.
-tracking, point correspondence, etc) to ensure appearance or temporal consistency.
-Such overhead could lead to much higher time and memory costs.
-These limitations make them incapable of meeting the diverse user requirements, necessitating the establishment of separate video editing workflows for different tasks.
-In recognition of these limitations, a compelling need arises to conceptualize and develop a simple and general-purpose video editing solution.
-In this paper, we introduce a unified yet simple framework, AnyV2V, designed to address (possibly) any video editing tasks.
-Our primary insight is that any video editing task can be effectively decomposed into two pivotal stages:
+The development of deep generative models [23] has led to significant advancements in content creation and manipulation, especially in digital images [43, 36, 4, 28, 9, 30].
+However, video generation and editing have not reached the same level of advancement as images [51, 6, 7, 24].
+In the context of video editing, training a large-scale video editing model presents considerable challenges due to the scarcity of paired data and the substantial computational resources required.
+To overcome these challenges, researchers proposed various approaches [16, 12, 52, 53, 32, 31, 19, 12, 60, 39, 5, 56, 27, 20], which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models.
+The zero-shot methods [16, 12, 27, 60] usually suffer from flickering issues due to a lack of temporal understanding.
+On the other hand, fine-tuning methods [52, 8, 53, 19, 20] require more time and computational overhead to edit videos.
+Moreover, all the methods can only adhere to certain types of edits.
+For example, a user might want to perform edits on visuals that are out-of-distribution from the learned text encoder (e.g.
+change a person to a character from their artwork).
+Thus, a highly customizable solution is more desired in video editing applications.
+It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ originality and creativity while producing high-quality results.
+With the above motivations, we aim to develop a video editing framework that requires no fine-tuning and caters to user demand.
+In this work, we present AnyV2V, designed to enhance the controllability of zero-shot video editing by decomposing the task into two pivotal stages:
   1. 1.
-    Apply an off-the-shelf task-specific image editing model over the first frame.
+    Apply image edit on the first frame with any off-the-shelf image editing model or human effort.
   2. 2.
-    Apply an image-to-video model to perform DDIM inversion on the source video and then inject the intermediate features to generate the new video.
-In the first stage, AnyV2V employs an existing task-specific image editing method to edit the first frame(Section 4.1).
-In the second stage, we aim to propagate the edited effect across the entire video while ensuring alignment with the source video.
-To achieve this, we follow an invert-and-generate framework [56, 38].
-In contrast to prior methods that employed text-to-video models [3, 4] for DDIM inversion [47], we employ the image-to-video models [62, 12, 44] for DDIM inversion(Section 4.2) to enable the first-frame condition.
-With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model is able to generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video.
-To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in convolution layers, spatial attention layers and temporal attention layers in the denoising U-Net of the I2V model.
-These techniques are discussed in detail in Section 4.3 and Section 4.4, and we demonstrate their effectiveness by conducting an ablation study in Section 5.5.
-By performing the two-stage editing process, AnyV2V effectively offloads the editing operation to existing image editing tools.
-This disentanglement helps AnyV2V excel in:
-  1. 1.
-    Compatibility: AnyV2V has superior compatibility with all the image editing methods.
-We demonstrate that AnyV2V can seamlessly build on top of advanced image editing methods, such as InstructPix2Pix [6], InstantID [51], NST [18], AnyDoor [11] to perform diverse types of editing.
-  2. 2.
-    Simplicity: AnyV2V is a tuning-free approach without requiring any additional video features to achieve high appearance and temporal consistency.
-In this paper, we showcase AnyV2V’s versatility and effectiveness by the qualitative and quantitative results.
-We comprehensively evaluate our method on four important tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing and (4) identity manipulation.
-We collected a new dataset of roughly 100 examples based on the four tasks to evaluate the performance of AnyV2V and other baselines [20, 15, 55].
-We show that AnyV2V is among the first to perform (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation in the video domain.
-Moreover, the AnyV2V exhibits 25% human preference improvement over the previous best model [20] in (1) prompt-based editing.
-We also perform a comprehensive ablation study to show the impact of our design choices.
-Our contributions are summarized as follows:
-  1. 1.
-    We introduce AnyV2V, a plug-and-play unified framework tailored for a diverse range of video-to-video editing tasks.
-  2. 2.
-    To the best of our knowledge, we are the first to perform video editing using pre-trained I2V models, marking a novel paradigm in this field.
-  3. 3.
-    Universal compatibility with image editing methods. AnyV2V is compatible with any image editing method, extending any image editing method into the video domain at no cost.
-  4. 4.
-    We show, both quantitatively and qualitatively, that our method outperforms existing state-of-the-art baselines on prompt-based editing and exhibits robust performance on three novel tasks: reference-based style transfer, subject-driven editing, and identity manipulation.
-2 Related Work
-2.0.1 Text-to-Video (T2V) Generation.
-Recent advancements in diffusion models [27, 46] have led to a huge development in text-to-video (T2V) generation.
-Existing diffusion-based T2V models generally follow the paradigm of extending text-to-image (T2I) generation models to the video domain, where additional temporal modelling layers are added to the T2I model to enable video generation.
-Recent T2V generation models are developed from pixel-space diffusion models [29, 26, 19], latent diffusion models (LDMs) [63, 1, 5, 24, 8, 53], as well as diffusion transformers [25, 10, 41].
-LDM-based [45] T2V generation models are the most extensively studied methods due to their high efficiency and the open-sourced Stable Diffusion [45] models.
-T2V models have been shown to be foundational models for a wide range of applications such as I2V generation [57], personalized video generation [39] and video editing [3].
-2.0.2 Image-to-Video (I2V) Generation.
-To achieve more precise control over the video generation process, a series of I2V generation methods employing additional image conditions in the video generation process have emerged.
-Earlier efforts such as VideoComposer [52] and VideoCrafter1 [8] use the input image as a style reference and cannot fully preserve the visual details in the output videos.
-More recently, methods such as DynamiCrafter [57], I2VGen-XL [62], SEINE [12] and ConsistI2V [44] have overcome this problem and can generate videos that are more consistent with the input image (often used as the first frame of the video).
-I2V generation models have also been employed in a wide range of applications such as long video generation [12, 59, 44], generative frame interpolation [57, 12] and video storytelling [57, 59].
-However, to the best of our knowledge, no methods have applied I2V generation models for video editing.
-We aim to investigate this area and propose a highly flexible training-free framework for video editing using I2V generation models.
-2.0.3 Semantic Image and Video Editing.
-Visual generation has attracted considerable attention within the field; however, visual manipulation also represents a significant and popular area of interest.
-The research area of image editing encompasses a wide range of tasks.
-The popular tasks include stylization [18, 13, 37], identity manipulation [51], subject-driven editing [34, 22, 11], and localized editing [40, 2, 48, 33].
-One interesting direction is to train a model to understand text inputs to perform general image editing tasks [60, 6, 49].
-However, image editing remained an open and challenging problem as none of the work was able to tackle all the tasks effectively [31].
-On the other hand, the video editing problem was often perceived as an extension of image editing problems, incorporating the time dimension.
-Recent video editing works [20, 54, 42, 15, 58, 7] employ text prompts to execute video editing tasks.
-These approaches have opened new avenues for interactive and intuitive video manipulation, allowing users to guide the editing process through natural language.
-However, such methodologies fail to offer precise control to the user, often resulting in edits that may not fully align with the user’s intentions or the desired level of detail due to the ambiguity of natural language and the limit of model capability.
-Our work is the first to enable precise video editing using pre-trained Image-to-Video (I2V) models.
-More importantly, our paradigm merges the two problems into a single domain to create a unified solution, offering high controllability in video editing tasks.
-3 Preliminary
-3.0.1 Image-to-Video (I2V) Generation Models.
-In this work, we focus on leveraging latent diffusion-based [45] I2V generation models for video editing.
-Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$.
-The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
-$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (1)
-$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V,$ (2)
-where $z$ is the input hidden state to the self-attention layer and $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable projection matrices that map $z$ onto query, key and value vectors, respectively.
-For spatial self-attention, $z$ represents a sequence of spatial tokens from each frame.
-For temporal self-attention, $z$ is composed of tokens located at the same spatial position across all frames.
-3.0.2 DDIM Inversion.
-The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM [47] sampling algorithm.
-The reverse process of DDIM sampling, known as DDIM inversion [38, 16], allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
-3.0.3 Plug-and-Play (PnP) Diffusion Features.
-Tumanyan et al.
-[49] proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g.
-legs or torso of a human body) during the image generation process.
-Given an input source image $I^{S}$ and a target prompt $P$, PnP first performs DDIM inversion to obtain the image’s corresponding noise $\{\mathbf{z}^{S}_{t}\}_{t=1}^{T}$ at each time step $t$.
-It then collects the convolution features $\{f^{l}_{t}\}$ and attention scores $\{A^{l}_{t}\}$ from some predefined layers $l$ at each time step $t$ of the backward diffusion process $\mathbf{z}_{t-1}^{S}=\epsilon_{\theta}(\mathbf{z}_{t}^{S},\varnothing,t)$, where $\varnothing$ denotes the null text prompt during denoising.
-To generate the edited image $I^{*}$, PnP starts from the initial noise of the source image (i.e.
-$\mathbf{z}_{T}^{*}=\mathbf{z}_{T}^{S}$) and performs feature injection during denoising: $\mathbf{z}_{t-1}^{*}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},P,t,\{f^{l}_{t},A^{l}_{t}\})$, where $\epsilon_{\theta}(\cdot,\cdot,\cdot,\{f^{l}_{t},A^{l}_{t}\})$ represents the operation of replacing the intermediate feature and attention scores $\{f^{l*}_{t},A^{l*}_{t}\}$ with $\{f^{l}_{t},A^{l}_{t}\}$.
-This feature injection mechanism ensures $I^{*}$ to preserve the layout and structure from $I^{S}$ while reflecting the description in $P$.
-To control the feature injection strength, PnP also employs two thresholds $\tau_{f}$ and $\tau_{A}$ such that the feature and attention scores are only injected in the first $\tau_{f}$ and $\tau_{A}$ denoising steps.
-Our method extends this feature injection mechanism to I2V generation models, where we inject features in convolution, spatial, and temporal attention layers.
-We show the detailed design of AnyV2V in Section 4.
+    Leverage the innate knowledge of the image-to-video (I2V) model to generate the edited video with the edited first frame, source video latent, and the intermediate temporal features.
+Our objective is to propagate the edited first frame across the entire video while ensuring alignment with the source video.
+To achieve this, we employ I2V models [59, 10, 42] for DDIM inversion to enable first-frame conditioning.
+With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model can generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video.
+To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in the I2V model.
+The two-stage editing process effectively offloads the editing operation to existing image editing tools.
+This design (detail in Section 3) helps AnyV2V excel in:
+  • Compatibility: It provides a highly customized interface for a user to perform video edits on any modality. It can seamlessly integrate any image editing methods to perform diverse editing.
+  • Simplicity: It does not require any fine-tuning nor additional video features like previous works [19, 53, 38] to achieve high appearance and temporal consistency for video editing tasks.
+From our findings, without any fine-tuning, AnyV2V can perform video editing tasks beyond the scope of current publicly available methods, such as reference-based style transfer, subject-driven editing, and identity manipulation.
+AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models [16, 12, 53].
+We show both quantitatively and qualitatively that our method outperforms existing SOTA baselines in Section 4.2 and Appendix B.
+AnyV2V is favoured in 69.7% of samples for prompt alignment and 46.2% overall preference in human evaluation, while the best baseline only achieves 31.7% and 20.7%, respectively.
+Our method also reaches the highest CLIP-Text score of 0.2932 in text alignment and a competitive CLIP-Image score of 0.9652 in temporal consistency.
+All these achievements are thanks to the AnyV2V’s design to harness the power of off-the-shelf image editing models from advanced image editing research.
+Through a comprehensive study and evaluation of the effectiveness of our design, our key observation is that the inverted noise latent and feature injection serve as critical components to guide the video motion, and the I2V model itself has good capabilities in generating motions.
+We also found that by inverting long videos exceeding the I2V models’ training frames, the inverted latents enable the I2V model to produce longer videos, making long video editing possible.
+To summarize, The main contributions of our work are three-fold:
+  • We proposed AnyV2V as a first fundamentally different solution for video editing, treating video editing as a simpler image editing problem.
+  • We showed that AnyV2V can support long video editing by inverting videos that extend beyond the training frame lengths of I2V models.
+  • Our extensive experiments showcased the superior performance of AnyV2V when compared to the existing SOTA methods, highlighting the potential of leveraging I2V models for video editing.
+2 Related Works
+Video generation has attracted considerable attention within the field [6, 7, 37, 51, 25, 57, 21, 50, 55, 10, 3, 42, 59].
+However, video manipulation also represents a significant and popular area of interest.
+Initial attempts, such as Tune-A-Video [53] and VideoP2P [33] involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion.
+The concurrent works at that time such as Pix2Video [5] and Fate-Zero [39] go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information.
+The progressively propagates to other frames edits.
+Subsequent developments have enhanced the results but generally follow the two paradigms  [16, 52, 12, 56, 5, 38, 20, 19, 14, 8, 27, 60, 11].
+Control-A-Video [8] and ControlVideo [60] leveraged ControlNet [58] for extra spatial guidance.
+TokenFlow [16] leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit.
+Fairy [52] followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits.
+VideoSwap [19] requires additional parameter tuning and video feature extraction (e.g.
+tracking, point correspondence, etc) to ensure appearance and temporal consistency.
+CoDeF [38] allows the first image edit to propagate the other frames with one-shot tuning.
+UniEdit [2] leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model [51].
+Table 1: Comparison with different video editing methods and the type of editing tasks.
+Method
+Prompt-based
+Editing
+Reference-based
+Style Transfer
+Subject-Driven
+Editing
+Identity
+Manipulation
+Tuning-Free? Backbone
+Tune-A-Video [53] Stable Diffusion [43]
+Pix2Video [5] SD-Depth  [43]
+Gen-1 [14] Stable Diffusion [43]
+TokenFlow [16] Stable Diffusion [43]
+FLATTEN [12] Stable Diffusion [43]
+Fairy [52] Stable Diffusion [43]
+ControlVideo [60] ControlNet [58]
+CoDeF [38] ControlNet [58]
+VideoSwap [19] AnimateDiff [20]
+UniEdit [2] Any T2V [51]
+AnyV2V (Ours) Any I2V [42, 59, 10]
+However, none of the methods can offer precise control to users, as the edits may not align with the user’s exact intentions or desired level of detail, often due to the ambiguity of natural language and the constraints of the model’s capabilities.
+For example, VideoP2P [32] is restricted to only word-swapping prompts due to the reliance on cross-attention.
+There is a clear need for a more precise and comprehensive solution for video editing tasks.
+Our work AnyV2V is the first work to empower a diverse array of video editing tasks.
+We compare AnyV2V with the existing methods in Table 1.
+As can be seen, our method excels in its applicability and compatibility.
 3 AnyV2V
-Figure 2: AnyV2V framework.
-Our framework takes a source video $V^{S}$ as input.
+Figure 2: AnyV2V takes a source video $V^{S}$ as input.
 In the first stage, we apply a block-box image editing method on the first frame $I_{1}$ according to the editing task.
 In the second stage, the source video is inverted to initial noise $z_{T}^{S}$, which is then denoised using DDIM sampling.
-During the sampling process, we extract spatial features, spatial attention, and temporal attention from the image-to-video models’ decoder layers.
-To generate our edited video, we perform a DDIM sampling by fixing $z_{T}^{*}$ as $z_{T}^{T}$ and use the edited first frame as the conditional signal.
-During the sampling, we inject the features and attention into corresponding layers of the model.
+During the sampling process, we extract spatial convolution, spatial attention, and temporal attention features from the I2V models’ decoder layers.
+To generate the edited video, we perform a DDIM sampling by fixing $z_{T}^{*}$ as $z_{T}^{T}$ and use the edited first frame as the conditional signal.
+During sampling, we inject the features and attention into corresponding layers of the model.
 Our method presents a two-stage approach to video editing.
-Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ dentes the video length.
-We extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$, where $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc.
-In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$ and the semantic information represented in the edited first frame $I^{*}_{1}$ and the target prompt $\mathbf{s^{*}}$.
+Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ denotes the video length, we first extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$.
+$C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc.
+In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$, the semantic information represented in the edited first frame $I^{*}_{1}$, and the target prompt $\mathbf{s^{*}}$.
 An overall illustration of our video editing pipeline is shown in Figure 2.
+We provide a discussion of the preliminaries of our method in Appendix C.
 In this section, we explain each core component of our method.
@@
 In visual manipulation, controllability is a key element in performing precise editing.
-Current video editing models lack this controllability as most of the existing methods can only edit videos based on target text prompts (e.g.
-TokenFlow [20]) or text instructions (e.g.
-InsV2V [14]).
-Our AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame.
-This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [18, 21, 37], mask-based image editing [40, 2], image inpainting [48, 33], identity-preserving image editing [51], and subject-driven image editing [11, 34, 22].
-This stage is highly flexible and it can even be done by human experts.
+AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame.
+This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [15, 17, 34, 48], mask-based image editing [36, 1], image inpainting [45, 29], identity-preserving image editing [49], and subject-driven image editing [9, 30, 18].
 3.2 Structural Guidance using DDIM Inverison
@@
 3.3 Appearance Guidance via Spatial Feature Injection
-Our empirical observation (cf.
-Section 5.5) suggests that I2V generation models already have some editing capabilities by only using the edited first frame and DDIM inverted noise as the model input.
+Our empirical observation (Section 4.3) suggests that I2V generation models already have some editing capabilities by only using the edited first frame and DDIM inverted noise as the model input.
 However, we find that this simple approach is often unable to correctly preserve the background in the edited first frame and the motion in the source video, as the conditional signal from the source video encoded in the inverted noise is limited.
@@
 During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$.
-We preserve two types of hidden features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers.
+We preserve two types of features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers.
 We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors.
@@
 We also only apply temporal attention injection in the first $\tau_{ta}$ steps during sampling.
-Overall, combining the spatial and temporal feature injection mechanisms, the denoising process of our AnyV2V can be represented by:
-$\mathbf{z}^{*}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},I^{*},\mathbf{s}^{*},t,\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\}).$ (4)
+3.5 Putting it Together
+Overall, combining the spatial and temporal feature injection mechanisms, we replace the editing branch features $\{f^{*l_{1}},Q_{s}^{*l_{2}},K_{s}^{*l_{2}},Q_{t}^{*l_{3}},K_{t}^{*l_{3}}\}$ with the features from the source denoising branch:
+$\mathbf{z}^{*}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},I^{*},\mathbf{s}^{*},t\ ;\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\}),$ (2)
+where $\epsilon_{\theta}(\cdot\ ;\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\})$ denotes the feature replacement operation across different layers $l_{1},l_{2},l_{3}$.
 Our proposed spatial and temporal feature injection scheme enables tuning-free adaptation of I2V generation models for video editing.
@@
 4 Experiments
-In this section, we first introduce four video editing tasks that we focus on evaluating our AnyV2V framework: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation.
-We then assess the effectiveness of AnyV2V both quantitatively and qualitatively on these four tasks.
-For (1) prompt-based editing, we compare AnyV2V with SoTA text-guided video editing baselines.
-For the three novel tasks (2)(3)(4), as these tasks require reference images to guide video editing, current text-based video editing methods are incapable of performing such edits.
-We thus compare the performance of AnyV2V using different I2V generation backbones.
-5.1 Tasks Definition
+Figure 3: AnyV2V is robust in a wide range of prompt-based editing tasks while preserving the background. The results align the most with the text prompt and maintain high motion consistency.
+Figure 4: With different image editing models, AnyV2V can achieve a wide range of editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation.
+4.1 Implementation Details
+We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl.
+[59], ConsistI2V [42] and SEINE [10].
+For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps.
+We use the DDIM [44] sampler and set $T$ to the default values of the selected I2V models.
+Following PnP [47], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections.
+During sampling, we apply text classifier-free guidance (CFG) [22] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits.
+To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [4], style transfer model Neural Style Transfer (NST) [15], subject-driven image editing model AnyDoor [9], and identity-driven image editing model InstantID [49].
+We experiment with only the successfully edited frames, which is crucial for our method.
+We conducted all the experiments on a single Nvidia A6000 GPU.
+To edit a 16-frame video, it requires around 15G GPU memory and around 100 seconds for the whole inference process.
+We refer readers to Appendix A for more discussions on our implementation details and hyperparameter settings.
+[…]
+Method Human Evaluation $\uparrow$ CLIP Scores $\uparrow$
+Alignment Overall CLIP-Text CLIP-Image
+Tune-A-Video [53] 15.2% 2.1% 0.2902 0.9704
+[…]
+[…]
+[…]
+[…]
+AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
+4.2 Evaluation Results
+As shown in Figure 3 and 4, AnyV2V can perform the following video editing tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation.
+we compare AnyV2V against three baseline models Tune-A-Video [53], TokenFlow [16] and FLATTEN [12] for (1) prompt-based editing.
+Since there exists no publicly available baseline method for task (2) (3) (4), we evaluate the performance of three I2V generation models under AnyV2V.
+We included a more comprehensive evaluation in Appendix B.
+Prompt-based Editing Unlike the baseline methods [53, 16, 12] that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged.
+Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix [4], AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue.
+Additionally, it maintains the original video’s background and fidelity, whereas, in comparison, baseline methods often alter the color tone and shape of objects, as illustrated in Figure 3.
+Also, for motion tasks such as adding snowing weather, I2V models from AnyV2V provide inherent support for animating the snowing while the baseline methods would result in flickering.
+For quantitative evaluations, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting, and also compute the metrics CLIP-Text for text alignment and CLIP-Image for temporal consistency.
+In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model [41] and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames.
+Table 2 shows that AnyV2V generally achieves high text-alignment and temporal consistency, while AnyV2V with I2VGen-XL backbone is the most preferred method because it does not edit the video precisely.
+Style Transfer, Subject-Driven Editing and Identity Manipulation For these novel tasks, we stress the alignment with the reference image instead of the text prompt.
+As shown in Figure 4, we can observe that for task (2), AnyV2V can capture one particular style that is tailor-made, and even if the style is not learned by the text encoder.
+In the examples, AnyV2V captures the style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately.
+For task (3), AnyV2V can replace the subject in the video with other subjects even if the new subject differs from the original subject.
+In the examples, a cat is replaced with a dog according to the reference image and maintains highly aligned motion and background as reflected in the source video.
+A car is replaced by our desired car while the wheel is still spinning in the edited video.
+For task (4), AnyV2V can swap a person’s identity to anyone.
+We report both human evaluation results and find that AnyV2V with I2VGen-XL backbone is the most preferred method in terms of reference alignment and overall performance, plotted in Table 5, which is in Appendix B.
+I2V Backbones We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively.
+It has a good generalization ability to produce consistent motions in the video with high visual quality.
+AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its training data, thus harming the visual quality.
+AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
+Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models [10, 42, 59] are mostly trained on video data that contains only 16 frames.
+To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V [42] and SEINE [10].
+However, we find that such an experiment setup cannot maintain semantic consistency in our case.
+As many works in extending video generation exploit the initial latent to generate longer video [40, 54], we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output.
+Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
+Figure 5: AnyV2V can edit video length beyond the training frame while maintaining motion consistency.
+The first row is the source video frames while the second rows are the edited.
+The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix [4].
+4.3 Ablation Study
+To verify the effectiveness of our design choices, we conduct an ablation study by iteratively disabling the three core components in our model: temporal feature injection, spatial feature injection, and DDIM inverted latent as initial noise.
+We use AnyV2V (I2VGen-XL) and a subset of 20 samples in this ablation study and report both the frame-wise consistency results using CLIP-Image score in Table 3 and qualitative comparisons in Figure 6.
+We provide more ablation analysis of other design considerations of our model in the Appendix.
+Figure 6: Visual comparisons of AnyV2V’s editing results after disabling temporal feature injection (T.I.), spatial feature injection (S.I.) and DDIM inverted initial noise (D.I.).
+[…]
+Effectiveness of Spatial Feature Injection As shown in Table 3, we observe a drop in the CLIP-Image score after removing the spatial feature injection mechanisms from our model, indicating that the edited videos are not smoothly progressed across consecutive frames and contain more appearance and motion inconsistencies.
+Further illustrated in the third row of Figure 6, removing spatial feature injection will often result in incorrect subject appearance and pose (as shown in the “ballet dancing” case) and degenerated background appearance (evident in the “couple sitting” case).
+These observations demonstrate that directly generating edited videos from the DDIM inverted noise is often not enough to fully preserve the source video structures, and the spatial feature injection mechanisms are crucial for achieving better editing results.
+DDIM Inverted Noise as Structural Guidance Finally, we observe a further decrease in CLIP-Image scores and a significantly degraded visual appearance in both examples in Figure 6 after replacing the initial DDIM inverted noise with random noise during sampling.
+This indicates that the I2V generation models become less capable of animating the input image when the editing prompt is completely out-of-domain and highlights the importance of the DDIM inverted noise as the structural guidance of the edited videos.
+Table 3: Ablation study results for AnyV2V (I2VGen-XL). T. Injection and S. Injection correspond to temporal and spatial feature injection mechanisms, respectively.
+Model CLIP-Image $\uparrow$
+AnyV2V (I2VGen-XL) 0.9648
+AnyV2V (I2VGen-XL) w/o T. Injection 0.9652
+AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection 0.9637
+AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection & DDIM Inversion 0.9607
+5 Conclusion
+In this paper, we presented AnyV2V, a novel unified framework for video editing.
+Our framework is training-free, highly cost-effective, and can be applied to any image editing model and I2V generation model.
+To perform video editing with high precision, we propose a two-stage approach to first edit the initial frame of the source video and then condition an I2V model with the edited first frame and the source video features and inverted latents to produce the edited video at any length.
+Comprehensive experiments have shown that our method achieves outstanding outcomes across a broad spectrum of applications that are beyond the scope of existing SOTA methods while achieving superior results on both common video metrics and human evaluation.
+For future work, we aim to find a tuning-free method to bridge the I2V properties into T2V models so that we can leverage existing strong T2V models for video editing.
+Appendix
+Appendix A Discussion on Model Implementation Details
+When adapting our AnyV2V to various I2V generation models, we identify two sets of hyperparameters that are crucial to the final video editing results.
+They are (1) selection of U-Net decoder layers ($l_{1}$, $l_{2}$ and $l_{3}$) to perform convolution, spatial attention and temporal attention injection and (2) Injection thresholds $\tau_{conv}$, $\tau_{sa}$ and $\tau_{ta}$ that control feature injections to happen in specific diffusion steps.
+In this section, we provide more discussions and analysis on the selection of these hyperparameters.
+A.1 U-Net Layers for Feature Injection
+Figure 7: Visualizations of the convolution, spatial attention and temporal attention features during video sampling for I2V generation models’ decoder layers.
+We feed in the DDIM inverted noise to the I2V models such that the generated videos (first row) are reconstructions of the source video.
+To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL [59], ConsistI2V [42] and SEINE [10].
+Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e.
+average attention weights for all other tokens attending to the current token).
+The results are shown in Figure 7.
+According to the figure, we observe that the intermediate convolution features from different I2V models show similar characteristics during video generation: earlier layers in the U-Net decoder produce features that represent the overall layout of the video frames and deeper layers capture the high-frequency details such as edges and textures.
+We choose to set $l_{1}=4$ for convolution feature injection to inject background and layout guidance to the edited video without introducing too many high-frequency details.
+For spatial and temporal attention scores, we observe that the spatial attention maps tend to represent the semantic regions in the video frames while the temporal attention maps highlight the foreground moving subjects (e.g.
+the running woman in Figure 7).
+One interesting observation for I2VGen-XL is that its spatial attention operations in deeper layers almost become hard attention, as the spatial tokens only attend to a single or very few tokens in each frame.
+We propose to inject features in decoder layers 4 to 11 ($l_{2}=l_{3}=\{4,5,...,11\}$) to preserve the semantic and motion information from the source video.
+A.2 Ablation Analysis on Feature Injection Thresholds
+We perform additional ablation analysis using different feature injection thresholds to study how these hyperparameters affect the edited video.
+Effect of Spatial Injection Thresholds $\tau_{conv},\tau_{sa}$
+We study the effect of disabling spatial feature injection or using different $\tau_{conv}$ and $\tau_{sa}$ values during video editing and show the qualitative results in Figure 8.
+We find that when spatial feature injection is disabled, the edited videos fail to fully adhere to the layout and motion from the source video.
+When spatial feature injection thresholds are too high, the edited videos are corrupted by the high-frequency details from the source video (e.g.
+textures from the woman’s down jacket in Figure 8).
+Setting $\tau_{conv}=\tau_{sa}=0.2T$ achieves a desired editing outcome for our experiments.
+Figure 8: Hyperparameter study on spatial feature injection.
+We find that $\tau_{sa}=0.2T$ is the best setting for maintaining the layout and structure in the edited video while not introducing unnecessary visual details from the source video.
+$\tau_{c,s}$ represents $\tau_{conv}$ and $\tau_{sa}$.
+(Editing prompt: teddy bear running.
+The experiment was conducted with the I2Vgen-xl backbone.
+Effect of Temporal Injection Threshold $\tau_{ta}$
+We study the hyperparameter of temporal feature injection threshold $\tau_{ta}$ in different settings and show the results in Figure 9.
+We observe that in circumstances where $\tau_{ta}<0.5T$ ($T$ is the total denoising steps), the motion guidance is too weak that it leads to only partly aligned motion with the source video, even though the motion itself is logical and smooth.
+At $\tau_{ta}>0.5T$, the generated video shows a stronger adherence to the motion but distortion occurs.
+We employ $\tau_{ta}=0.5T$ in our experiments and find that this value strikes the perfect balance on motion alignment, motion consistency, and video fidelity.
+Figure 9: Hyperparameter study on temporal feature injection.
+We find that $\tau_{ta}=0.5T$ to be the optimal setting as it balances motion alignment, motion consistency, and video fidelity.
+(Editing prompt: darth vader walking.
+The experiment was conducted with the SEINE backbone.
+Appendix B Evaluation
+B.1 Tasks Definition
   • 1.
@@
     Identity Manipulation: Identity manipulation allows the user to manipulate video content by replacing a person with another person’s identity in the video based on an input image of the target person.
-5.2 Implementation Details
-We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl.
-[62], ConsistI2V [44] and SEINE [12].
-For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps.
-We use the DDIM [47] sampler and set $T$ to the default values of the selected I2V models.
-Following PnP [50], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections.
-During sampling, we apply text classifier-free guidance (CFG) [28] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits.
-We refer readers to the supplementary materials for more discussions on our implementation details and hyperparameter settings.
-To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [6], style transfer model Neural Style Transfer (NST) [18], subject-driven image editing model AnyDoor [11], and identity-driven image editing model InstantID [51].
-We experiment with only the successfully edited frames, which is crucial for our method.
-5.3 Quantitative Evaluations
-5.3.1 Prompt-based Editing.
+B.2 Quantitative Evaluations
+Prompt-based Editing
 For (1) prompt-based editing, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting.
@@
 Our automatic evaluations in Table 2 demonstrate that our model is competitive in prompt-based editing compared to baseline methods.
-5.3.2 Reference-based Style Transfer; Identity Manipulation and Subject-driven Editing
+Reference-based Style Transfer; Identity Manipulation and Subject-driven Editing
 For novel tasks (2), (3) and (4), we evaluate the performance of three I2V generation models using human evaluations and show the results in Table 5.
@@
 Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing [28], we also report the image editing success rate in Table 5 to clarify that our method relies on a good image frame edit.
-[…]
-Task Prompt-based Editing
-Method Human Evaluation $\uparrow$ CLIP Scores $\uparrow$
-Alignment Overall CLIP-Text CLIP-Image
-Tune-A-Video [55] 15.2% 2.1% 0.2902 0.9704
-[…]
-[…]
-[…]
-[…]
-AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
-[…]
-Task Reference-based Subject-driven Identity
-Style Transfer Editing Manipulation
-Image Editing Method NST [18] AnyDoor [11] InstantID [51]
-Image Editing Success Rate $\approx$90$\approx$10$\approx$80%
-Human Evaluation Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$
-[…]
-[…]
-AnyV2V (I2VGen-XL) 100.0% 76.9% 93.9% 84.8% 90.1% 45.4%
-5.4 Qualitative Results
-5.4.1 Prompt-based Editing.
+B.3 Qualitative Results
+Prompt-based Editing
 By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing.
@@
 It is also worth mentioning that our approach is far more consistent on some motion tasks such as adding snowing weather, due to the I2V model’s inherent support for animating still scenes.
-The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modelling.
-Figure 3: AnyV2V is robust in a wide range of prompt-based editing tasks while maintaining the background.
-The generated results align the most with the text prompt and also maintain high motion consistency.
-Figure 4: With different image editing models, AnyV2V can achieve a wide range of editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation.
-5.4.2 Reference-based Style Transfer.
+The baseline methods, on the other hand, can add snow to individual frames but cannot generate the effect of snow falling, as the per-frame or one-shot editing methods lack the ability of temporal modeling.
+Reference-based Style Transfer
 Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST [15] to obtain the edited frame.
@@
 As demonstrated in Figure 4, our method captures the distinctive style of Vassily Kandinsky’s artwork “Composition VII” and Vincent Van Gogh’s artwork “Chateau in Auvers at Sunset” accurately, while such an edit is often hard to perform using existing text-guided video editing methods.
-5.4.3 Subject-driven Editing.
+Subject-driven Editing
 In our experiment, we employed a subject-driven image editing model AnyDoor [9] for the first frame editing.
@@
 In the second example, the car is replaced by our desired car while maintaining the rotation angle in the edited video.
-5.4.4 Identity Manipulation.
+Identity Manipulation
 By integrating the identity-preserved image personalization model InstantID [49] with ControlNet [58], this approach enables the replacement of an individual’s identity to create an initial frame.
@@
 It is possible to leverage other identity-preserved image personalization models and apply them to AnyV2V to preserve the background.
-Table 3: Ablation study results for AnyV2V (I2VGen-XL). T. Injection and S. Injection correspond to temporal and spatial feature injection mechanisms, respectively.
-Model CLIP-Image $\uparrow$
-AnyV2V (I2VGen-XL) 0.9648
-AnyV2V (I2VGen-XL) w/o T. Injection 0.9652
-AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection 0.9637
-AnyV2V (I2VGen-XL) w/o T. Injection & S. Injection & DDIM Inversion 0.9607
-5.5 Ablation Study
-To verify the effectiveness of our design choices, we conduct an ablation study by iteratively disabling the three core components in our model: temporal feature injection, spatial feature injection, and DDIM inverted latent as initial noise.
-We use AnyV2V (I2VGen-XL) and a subset of 20 samples in this ablation study and report both the frame-wise consistency results using CLIP-Image score in Table 3 and qualitative comparisons in Figure 5.
-We provide more ablation analysis of other design considerations of our model in the supplementary materials.
-Figure 5: Visual comparisons of AnyV2V’s editing results after disabling temporal feature injection (T.I.), spatial feature injection (S.I.) and DDIM inverted initial noise (D.I.).
-5.5.1 Effectiveness of Temporal Feature Injection.
-[…]
-5.5.2 Effectiveness of Spatial Feature Injection.
-As shown in Table 3, we observe a drop in the CLIP-Image score after removing the spatial feature injection mechanisms from our model, indicating that the edited videos are not smoothly progressed across consecutive frames and contain more appearance and motion inconsistencies.
-Further illustrated in the third row of Figure 5, removing spatial feature injection will often result in incorrect subject appearance and pose (as shown in the “ballet dancing” case) and degenerated background appearance (evident in the “couple sitting” case).
-These observations demonstrate that directly generating edited videos from the DDIM inverted noise is often not enough to fully preserve the source video structures, and the spatial feature injection mechanisms are crucial for achieving better editing results.
-5.5.3 DDIM Inverted Noise as Structural Guidance.
-Finally, we observe a further decrease in CLIP-Image scores and a significantly degraded visual appearance in both examples in Figure 5 after replacing the initial DDIM inverted noise with random noise during sampling.
-This indicates that the I2V generation models become less capable of animating the input image when the editing prompt is completely out-of-domain and highlights the importance of the DDIM inverted noise as the structural guidance of the edited videos.
-6 Conclusion
-In this paper, we present AnyV2V, a new unified framework for video editing.
-Our framework is training-free, highly cost-effective, and can be applied to any I2V model.
-To perform video editing with high precision, we propose a two-stage approach to first edit the initial frame of the source video and then condition an image-to-video (I2V) model with the initial frame.
-Comprehensive experiments have shown that our method achieves outstanding outcomes across a broad spectrum of applications and shows a high level of controllability.
-7 Limitations
-7.0.1 Inaccurate Edit from Image Editing Models.
-As our method relies on an initial frame edit, the image editing models are used.
-However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [32].
-For example, in the subject-driven video editing task, we found that AnyDoor [11] requires several tries to get a good editing result.
-Efforts are required in manually picking a good edited frame.
-We expect that in the future better image editing models will minimize such effort.
-7.0.2 Limited ability of I2V models.
-We found that the results from our method cannot follow the source video motion if the motion is fast (e.g.
-billiard balls hitting each other at full speed) or complex (e.g.
-a person clipping her hair).
-One possible reason is that the current popular I2V models are generally trained on slow-motion videos, such that lacking the ability to regenerate fast or complex motion even with motion guidance.
-We anticipate that the presence of a robust I2V model can address this issue.
-8 Potential Negative Impacts
-Misinformation spread and Privacy violations.
-As our technique allows for object manipulation, it can produce highly realistic yet completely fabricated videos of one individual or subject.
-There is a risk that harmful actors could exploit our system to generate counterfeit videos to disseminate false information.
-Moreover, the ability to create convincing counterfeit content featuring individuals without their permission undermines privacy protections, possibly leading to the illicit use of a person’s likeness for harmful purposes and damaging their reputation.
-These issues are similarly present in DeepFake technologies.
-To mitigate the risk of misuse, one proposed solution is the adoption of unseen watermarking, a method commonly used to tackle such concerns in image generation.
-Supplementary Material
-Appendix 0.A Discussion on Model Implementation Details
-When adapting our AnyV2V to various I2V generation models, we identify two sets of hyperparameters that are crucial to the final video editing results.
-They are (1) selection of U-Net decoder layers ($l_{1}$, $l_{2}$ and $l_{3}$) to perform convolution, spatial attention and temporal attention injection and (2) Injection thresholds $\tau_{conv}$, $\tau_{sa}$ and $\tau_{ta}$ that control feature injections to happen in specific diffusion steps.
-In this section, we provide more discussions and analysis on the selection of these hyperparameters.
-0.A.1 U-Net Layers for Feature Injection
-Figure 6: Visualizations of the convolution, spatial attention and temporal attention features during video sampling for I2V generation models’ decoder layers.
-We feed in the DDIM inverted noise to the I2V models such that the generated videos (first row) are reconstructions of the source video.
-To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL [62], ConsistI2V [44] and SEINE [12].
-Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e.
-average attention weights for all other tokens attending to the current token).
-The results are shown in Figure 6.
-According to the figure, we observe that the intermediate convolution features from different I2V models show similar characteristics during video generation: earlier layers in the U-Net decoder produce features that represent the overall layout of the video frames and deeper layers capture the high-frequency details such as edges and textures.
-We choose to set $l_{1}=4$ for convolution feature injection to inject background and layout guidance to the edited video without introducing too many high-frequency details.
-For spatial and temporal attention scores, we observe that the spatial attention maps tend to represent the semantic regions in the video frames while the temporal attention maps highlight the foreground moving subjects (e.g.
-the running woman in Figure 6).
-One interesting observation for I2VGen-XL is that its spatial attention operations in deeper layers almost become hard attention, as the spatial tokens only attend to a single or very few tokens in each frame.
-We propose to inject features in decoder layers 4 to 11 ($l_{2}=l_{3}=\{4,5,...,11\}$) to preserve the semantic and motion information from the source video.
-0.A.2 Ablation Analysis on Feature Injection Thresholds
-We perform additional ablation analysis using different feature injection thresholds to study how these hyperparameters affect the edited video.
-0.A.2.1 Effect of Spatial Injection Thresholds $\tau_{conv},\tau_{sa}$
-We study the effect of disabling spatial feature injection or using different $\tau_{conv}$ and $\tau_{sa}$ values during video editing and show the qualitative results in Figure 7.
-We find that when spatial feature injection is disabled, the edited videos fail to fully adhere to the layout and motion from the source video.
-When spatial feature injection thresholds are too high, the edited videos are corrupted by the high-frequency details from the source video (e.g.
-textures from the woman’s down jacket in Figure 7).
-Setting $\tau_{conv}=\tau_{sa}=0.2T$ achieves a desired editing outcome for our experiments.
-Source
-$\tau_{c,s}=0$
-$\tau_{c,s}=0.2T$
-$\tau_{c,s}=0.5T$
-$\tau_{c,s}=0.7T$
-$\tau_{c,s}=T$
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/source/015
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/0/015
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/2/015
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/5/015
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/7/015
-[width=]8Tables_and_Figures/Supp/Figures/video/spatial_inject/woman/10/015
-$\tau_{c,s}=0$
-$\tau_{c,s}=0.2T$
-$\tau_{c,s}=0.5T$
-$\tau_{c,s}=0.7T$
-$\tau_{c,s}=T$
-Figure 7: Hyperparameter study on spatial feature injection.
-We find that $\tau_{sa}=0.2T$ is the best setting for maintaining the layout and structure in the edited video while not introducing unnecessary visual details from the source video.
-$\tau_{c,s}$ represents $\tau_{conv}$ and $\tau_{sa}$.
-(Editing prompt: teddy bear running.
-Videos clickable in Adobe Acrobat.)
-0.A.2.2 Effect of Temporal Injection Threshold $\tau_{ta}$
-We study the hyperparameter of temporal feature injection threshold $\tau_{ta}$ in different settings and show the results in Figure 8.
-We observe that in circumstances where $\tau_{ta}<0.5T$ ($T$ is the total denoising steps), the motion guidance is too weak that it leads to only partly aligned motion with the source video, even though the motion itself is logical and smooth.
-At $\tau_{ta}>0.5T$, the generated video shows a stronger adherence to the motion but distortion occurs.
-We employ $\tau_{ta}=0.5T$ in our experiments and find that this value strikes the perfect balance on motion alignment, motion consistency, and video fidelity.
-Source
-$\tau_{ta}=0$
-$\tau_{ta}=0.2T$
-$\tau_{ta}=0.5T$
-$\tau_{ta}=0.7T$
-$\tau_{ta}=T$
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/src/frames/frame_015
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0_stp50/frames/frame_015
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.2_stp50/frames/frame_015
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.5_stp50/frames/frame_015
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp0.7_stp50/frames/frame_015
-[width=]8Tables_and_Figures/Supp/Figures/video/temporal_inject/seine/man_walking/cfg4_f0.2_spa0.2_cro0.0_tmp1.0_stp50/frames/frame_015
-$\tau_{ta}=0$
-$\tau_{ta}=0.2T$
-$\tau_{ta}=0.5T$
-$\tau_{ta}=0.7T$
-$\tau_{ta}=T$
-Figure 8: Hyperparameter study on temporal feature injection.
-We find that $\tau_{ta}=0.5T$ to be the optimal setting as it balances motion alignment, motion consistency, and video fidelity.
-(Editing prompt: darth vader walking.
-Videos clickable in Adobe Acrobat.)
-0.A.3 Backbones Performance
-Among the three backbones of our framework, We found that AnyV2V (I2VGen-XL) tends to be the most robust model both qualitatively and quantitatively.
-It has a good generalization ability to produce consistent motions in the video with high visual quality.
-AnyV2V (ConsistI2V) can generate consistent motion, but sometimes the watermark would appear due to its data used during the training stage, thus harming the visual quality.
-AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
-One interesting property of our framework is that although in some cases the motion does not fully align with the source video, it produces video with high motion consistency and reasonable motion that is highly similar to the source video, indicating the flexibility of how the video can be edited while maintaining the video fidelity.
-Appendix 0.B Human Evaluation Dataset
-Our human evaluation dataset contains in total of 89 samples.
+B.4 Human Evaluation
+B.4.1 Dataset
+Our human evaluation dataset contains a total of 89 samples that have been collected from https://www.pexels.com.
 For prompt-based editing, we employed InstructPix2Pix [4] to compose the examples.
@@
 Total 89
+B.4.2 Interface
+In the evaluation setup, we provide the evaluator with generated images from both the baseline models and the AnyV2V models.
+Evaluators are tasked with selecting videos that best align with the provided prompt or the reference image.
+Additionally, they are asked to express their overall preference for the edited videos.
+For a detailed view of the interface used in this process, please see Figure 10.
+Figure 10: The interface of individual evaluation.
+Appendix C Preliminary
+C.1 Image-to-Video (I2V) Generation Models
+In this work, we focus on leveraging latent diffusion-based [43] I2V generation models for video editing.
+Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$.
+The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
+$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (3)
+$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V,$ (4)
+where $z$ is the input hidden state to the self-attention layer and $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable projection matrices that map $z$ onto query, key and value vectors, respectively.
+For spatial self-attention, $z$ represents a sequence of spatial tokens from each frame.
+For temporal self-attention, $z$ is composed of tokens located at the same spatial position across all frames.
+C.2 DDIM Inversion
+The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM [44] sampling algorithm.
+The reverse process of DDIM sampling, known as DDIM inversion [35, 13], allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
+C.3 Plug-and-Play (PnP) Diffusion Features
+Tumanyan et al.
+[46] proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g.
+legs or torso of a human body) during the image generation process.
+Given an input source image $I^{S}$ and a target prompt $P$, PnP first performs DDIM inversion to obtain the image’s corresponding noise $\{\mathbf{z}^{S}_{t}\}_{t=1}^{T}$ at each time step $t$.
+It then collects the convolution features $\{f^{l}_{t}\}$ and attention scores $\{A^{l}_{t}\}$ from some predefined layers $l$ at each time step $t$ of the backward diffusion process $\mathbf{z}_{t-1}^{S}=\epsilon_{\theta}(\mathbf{z}_{t}^{S},\varnothing,t)$, where $\varnothing$ denotes the null text prompt during denoising.
+To generate the edited image $I^{*}$, PnP starts from the initial noise of the source image (i.e.
+$\mathbf{z}_{T}^{*}=\mathbf{z}_{T}^{S}$) and performs feature injection during denoising: $\mathbf{z}_{t-1}^{*}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},P,t,\{f^{l}_{t},A^{l}_{t}\})$, where $\epsilon_{\theta}(\cdot,\cdot,\cdot,\{f^{l}_{t},A^{l}_{t}\})$ represents the operation of replacing the intermediate feature and attention scores $\{f^{l*}_{t},A^{l*}_{t}\}$ with $\{f^{l}_{t},A^{l}_{t}\}$.
+This feature injection mechanism ensures $I^{*}$ to preserve the layout and structure from $I^{S}$ while reflecting the description in $P$.
+To control the feature injection strength, PnP also employs two thresholds $\tau_{f}$ and $\tau_{A}$ such that the feature and attention scores are only injected in the first $\tau_{f}$ and $\tau_{A}$ denoising steps.
+Our method extends this feature injection mechanism to I2V generation models, where we inject features in convolution, spatial, and temporal attention layers.
+We show the detailed design of AnyV2V in Section 3.
+[…]
+Task Reference-based Subject-driven Identity
+Style Transfer Editing Manipulation
+Image Editing Method NST [15] AnyDoor [9] InstantID [49]
+Image Editing Success Rate $\approx$90$\approx$10$\approx$80%
+Human Evaluation Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$ Align $\uparrow$ Overall $\uparrow$
+[…]
+[…]
+AnyV2V (I2VGen-XL) 100.0% 76.9% 93.9% 84.8% 90.1% 45.4%
+Appendix D Discussion
+D.1 Limitations
+Inaccurate Edit from Image Editing Models.
+As our method relies on an initial frame edit, the image editing models are used.
+However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [28].
+For example, in the subject-driven video editing task, we found that AnyDoor [9] requires several tries to get a good editing result.
+Efforts are required in manually picking a good edited frame.
+We expect that in the future better image editing models will minimize such effort.
+Limited ability of I2V models.
+We found that the results from our method cannot follow the source video motion if the motion is fast (e.g.
+billiard balls hitting each other at full speed) or complex (e.g.
+a person clipping her hair).
+One possible reason is that the current popular I2V models are generally trained on slow-motion videos, such that lacking the ability to regenerate fast or complex motion even with motion guidance.
+We anticipate that the presence of a robust I2V model can address this issue.
+D.2 License of Assets
+For image editing models, InstructPix2Pix [4] inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion.
+Neural Style Transfer [15] is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License.
+InstantID [49] is under Apache License 2.0.
+AnyDoor [9] is under the MIT License.
+For baselines, Tune-A-Video [53] is under Apache License 2.0, TokenFlow [16] is under MIT License andFLATTEN [12] is under Apache License 2.0.
+For the human evaluation dataset, the dataset has been collected from https://www.pexels.com, with all data governed by the terms outlined at https://www.pexels.com/license/.
+We decide to release AnyV2V code under the Creative Commons Attribution 4.0 License for easy access in the research community.
+D.3 Societal Impacts
+Postive Social Impact.
+AnyV2V has the potential to significantly enhance the capabilities of video editing systems, making it easier for a wider range of users to manipulate images.
+This could have numerous positive social impacts, as users would be able to achieve their editing goals without needing professional editing knowledge, such as using Photoshop or painting.
+Misinformation spread and Privacy violations.
+As our technique allows for object manipulation, it can produce highly realistic yet completely fabricated videos of one individual or subject.
+There is a risk that harmful actors could exploit our system to generate counterfeit videos to disseminate false information.
+Moreover, the ability to create convincing counterfeit content featuring individuals without their permission undermines privacy protections, possibly leading to the illicit use of a person’s likeness for harmful purposes and damaging their reputation.
+These issues are similarly present in DeepFake technologies.
+To mitigate the risk of misuse, one proposed solution is the adoption of unseen watermarking, a method commonly used to tackle such concerns in image generation.
+D.4 Safeguards
+It is crucial to implement proper safeguards and responsible AI frameworks when developing user-friendly video editing systems.
+For the human evaluation dataset, we manually collect a diverse range of images to ensure a balanced representation of objects from various domains.
+We only collect images that are considered safe.
+D.5 Ethical Concerns for Human Evaluation
+We believe our proposed human evaluation does not incur ethical concerns due to the following reasons: (1) the study does not involve any form of intervention or interaction that could affect the participants’ well-being.
+(2) Rating video content does not involve any physical or psychological risk, nor does it expose participants to sensitive or distressing material.
+(3) The data collected from participants will be entirely anonymous and will not contain any identifiable private information.
+(4) Participation in the study is entirely voluntary, and participants can withdraw at any time without any consequences.
+D.6 New Assets
+Our paper introduces several new assets including a human evaluation dataset and demo videos generated by AnyV2V.
+Each asset is thoroughly documented, detailing its creation, usage, and any relevant methodologies.
+The human evaluation dataset documentation includes details on how participant consent was obtained.
+The demo videos are provided as an anonymized zip file to comply with submission guidelines, with detailed instructions for use.
+All assets are shared under an open license to facilitate reuse and further research.
+Appendix E More AnyV2V Showcases
+Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix [4].
+Prompt used “Turn the couple into robots”, “Turn horse into zebra”, and “Turn the sand into snow”.
+Figure 12: The more recent model InstantStyle [48] can seamlessly plug in AnyV2V to perform reference-based style transfer video editing.
+The reference art image in the bottom left corner is used to retrieve the first image edit.
+Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor [9].
+The subject image in the bottom left corner is used to retrieve the first image edit.

### v4

@@
 AnyV2V can also support any video length.
-Our evaluation indicates that AnyV2V significantly outperforms other baseline methods in automatic and human evaluations by significant margin, maintaining visual consistency with the source video while achieving high-quality edits across all the editing tasks.
-Figure 1: AnyV2V is an evolving framework to handle all types of video-to-video editing tasks without any parameter tuning.
-AnyV2V disentangles video editing into two simpler problems: (1) Single image editing and (2) Image-to-video generation with video reference.
+Our evaluation shows that AnyV2V achieved CLIP-scores comparable to other baseline methods.
+Furthermore, AnyV2V significantly outperformed these baselines in human evaluations, demonstrating notable improvements in visual consistency with the source video while producing high-quality edits across all editing tasks.
+The code is available at https://github.com/TIGER-AI-Lab/AnyV2V.
 1 Introduction
-The development of deep generative models [23] has led to significant advancements in content creation and manipulation, especially in digital images [43, 36, 4, 28, 9, 30].
-However, video generation and editing have not reached the same level of advancement as images [51, 6, 7, 24].
+The development of deep generative models (Ho et al., 2020) has led to significant advancements in content creation and manipulation, especially in digital images (Rombach et al., 2022; Nichol et al., 2022; Brooks et al., 2023; Ku et al., 2024; Chen et al., 2023c; Li et al., 2023).
+However, video generation and editing have not reached the same level of advancement as images (Wang et al., 2023; Chen et al., 2023a; 2024; Ho et al., 2022).
 In the context of video editing, training a large-scale video editing model presents considerable challenges due to the scarcity of paired data and the substantial computational resources required.
-To overcome these challenges, researchers proposed various approaches [16, 12, 52, 53, 32, 31, 19, 12, 60, 39, 5, 56, 27, 20], which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models.
-The zero-shot methods [16, 12, 27, 60] usually suffer from flickering issues due to a lack of temporal understanding.
-On the other hand, fine-tuning methods [52, 8, 53, 19, 20] require more time and computational overhead to edit videos.
+To overcome these challenges, researchers proposed various approaches (Geyer et al., 2023; Cong et al., 2023; Wu et al., 2023a; b; Liu et al., 2023a; Liang et al., 2023; Gu et al., 2023b; Cong et al., 2023; Zhang et al., 2023d; Qi et al., 2023; Ceylan et al., 2023; Yang et al., 2023; Jeong & Ye, 2023; Guo et al., 2023), which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models.
+The zero-shot methods (Geyer et al., 2023; Cong et al., 2023; Jeong & Ye, 2023; Zhang et al., 2023d) usually suffer from flickering issues due to a lack of temporal understanding.
+On the other hand, fine-tuning methods (Wu et al., 2023a; Chen et al., 2023b; Wu et al., 2023b; Gu et al., 2023b; Guo et al., 2023) require more time and computational overhead to edit videos.
 Moreover, all the methods can only adhere to certain types of edits.
@@
 Thus, a highly customizable solution is more desired in video editing applications.
-It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ originality and creativity while producing high-quality results.
+It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ creativity while producing high-quality outputs.
+Figure 1: AnyV2V is an evolving framework to handle all types of video-to-video editing tasks without any parameter tuning.
+AnyV2V disentangles video editing into two simpler problems: (1) Single image editing and (2) Image-to-video generation with video reference.
 With the above motivations, we aim to develop a video editing framework that requires no fine-tuning and caters to user demand.
@@
 Our objective is to propagate the edited first frame across the entire video while ensuring alignment with the source video.
-To achieve this, we employ I2V models [59, 10, 42] for DDIM inversion to enable first-frame conditioning.
+To achieve this, we employ I2V models (Zhang et al., 2023c; Chen et al., 2023d; Ren et al., 2024) for DDIM inversion to enable first-frame conditioning.
 With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model can generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video.
@@
   • Compatibility: It provides a highly customized interface for a user to perform video edits on any modality. It can seamlessly integrate any image editing methods to perform diverse editing.
-  • Simplicity: It does not require any fine-tuning nor additional video features like previous works [19, 53, 38] to achieve high appearance and temporal consistency for video editing tasks.
+  • Simplicity: It does not require any fine-tuning nor additional video features like previous works (Gu et al., 2023b; Wu et al., 2023b; Ouyang et al., 2023) to achieve high appearance and temporal consistency for video editing tasks.
 From our findings, without any fine-tuning, AnyV2V can perform video editing tasks beyond the scope of current publicly available methods, such as reference-based style transfer, subject-driven editing, and identity manipulation.
-AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models [16, 12, 53].
+AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models (Geyer et al., 2023; Cong et al., 2023; Wu et al., 2023b).
 We show both quantitatively and qualitatively that our method outperforms existing SOTA baselines in Section 5.3 and Appendix B.
@@
 2 Related Works
-Video generation has attracted considerable attention within the field [6, 7, 37, 51, 25, 57, 21, 50, 55, 10, 3, 42, 59].
+Video generation has attracted considerable attention within the field (Chen et al., 2023a; 2024; OpenAI, ; Wang et al., 2023; Hong et al., 2022; Zhang et al., 2023a; Henschel et al., 2024; Wang et al., 2024c; Xing et al., 2023; Chen et al., 2023d; Bar-Tal et al., 2024; Ren et al., 2024; Zhang et al., 2023c).
 However, video manipulation also represents a significant and popular area of interest.
-Initial attempts, such as Tune-A-Video [53] and VideoP2P [33] involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion.
-The concurrent works at that time such as Pix2Video [5] and Fate-Zero [39] go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information.
+Initial attempts, such as Tune-A-Video (Wu et al., 2023b) and VideoP2P (Liu et al., 2023b) involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion.
+The concurrent works at that time such as Pix2Video (Ceylan et al., 2023) and Fate-Zero (Qi et al., 2023) go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information.
 The progressively propagates to other frames edits.
-Subsequent developments have enhanced the results but generally follow the two paradigms  [16, 52, 12, 56, 5, 38, 20, 19, 14, 8, 27, 60, 11].
-Control-A-Video [8] and ControlVideo [60] leveraged ControlNet [58] for extra spatial guidance.
-TokenFlow [16] leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit.
-Fairy [52] followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits.
-VideoSwap [19] requires additional parameter tuning and video feature extraction (e.g.
+Subsequent developments have enhanced the results but generally follow the two paradigms  (Geyer et al., 2023; Wu et al., 2023a; Cong et al., 2023; Yang et al., 2023; Ceylan et al., 2023; Ouyang et al., 2023; Guo et al., 2023; Gu et al., 2023b; Esser et al., 2023; Chen et al., 2023b; Jeong & Ye, 2023; Zhang et al., 2023d; Cheng et al., 2023).
+Control-A-Video (Chen et al., 2023b) and ControlVideo (Zhang et al., 2023d) leveraged ControlNet (Zhang et al., 2023b) for extra spatial guidance.
+TokenFlow (Geyer et al., 2023) leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit.
+Fairy (Wu et al., 2023a) followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits.
+VideoSwap (Gu et al., 2023b) requires additional parameter tuning and video feature extraction (e.g.
 tracking, point correspondence, etc) to ensure appearance and temporal consistency.
-CoDeF [38] allows the first image edit to propagate the other frames with one-shot tuning.
-UniEdit [2] leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model [51].
+CoDeF (Ouyang et al., 2023) allows the first image edit to propagate the other frames with one-shot tuning.
+UniEdit (Bai et al., 2024) leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model (Wang et al., 2023).
 Table 1: Comparison with different video editing methods and the type of editing tasks.
@@
 Tuning-Free? Backbone
-Tune-A-Video [53] Stable Diffusion [43]
-Pix2Video [5] SD-Depth  [43]
-Gen-1 [14] Stable Diffusion [43]
-TokenFlow [16] Stable Diffusion [43]
-FLATTEN [12] Stable Diffusion [43]
-Fairy [52] Stable Diffusion [43]
-ControlVideo [60] ControlNet [58]
-CoDeF [38] ControlNet [58]
-VideoSwap [19] AnimateDiff [20]
-UniEdit [2] Any T2V [51]
-AnyV2V (Ours) Any I2V [42, 59, 10]
+Tune-A-Video (Wu et al., 2023b) Stable Diffusion
+Pix2Video (Ceylan et al., 2023) SD-Depth
+Gen-1 (Esser et al., 2023) Stable Diffusion
+TokenFlow (Geyer et al., 2023) Stable Diffusion
+FLATTEN (Cong et al., 2023) Stable Diffusion
+Fairy (Wu et al., 2023a) Stable Diffusion
+ControlVideo (Zhang et al., 2023d) ControlNet
+CoDeF (Ouyang et al., 2023) ControlNet
+VideoSwap (Gu et al., 2023b) AnimateDiff
+UniEdit (Bai et al., 2024) Any T2V Models
+AnyV2V (Ours) Any I2V Models
 However, none of the methods can offer precise control to users, as the edits may not align with the user’s exact intentions or desired level of detail, often due to the ambiguity of natural language and the constraints of the model’s capabilities.
-For example, VideoP2P [32] is restricted to only word-swapping prompts due to the reliance on cross-attention.
+For example, VideoP2P (Liu et al., 2023a) is restricted to only word-swapping prompts due to the reliance on cross-attention.
 There is a clear need for a more precise and comprehensive solution for video editing tasks.
@@
 As can be seen, our method excels in its applicability and compatibility.
+3 Preliminary
+3.1 Image-to-Video (I2V) Generation Models
+In this work, we focus on leveraging latent diffusion-based (Rombach et al., 2022) I2V generation models for video editing.
+Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$.
+The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
+$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (1)
+$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V,$ (2)
+where $z$ is the input hidden state to the self-attention layer and $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable projection matrices that map $z$ onto query, key and value vectors, respectively.
+For spatial self-attention, $z$ represents a sequence of spatial tokens from each frame.
+For temporal self-attention, $z$ is composed of tokens located at the same spatial position across all frames.
+3.2 DDIM Inversion
+The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM (Song et al., 2020) sampling algorithm.
+The reverse process of DDIM sampling, known as DDIM inversion (Mokady et al., 2023; Dhariwal & Nichol, 2021), allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
+3.3 Plug-and-Play (PnP) Diffusion Features
+Tumanyan et al.
+(2023a) proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g.
+legs or torso of a human body) during the image generation process.
+Given an input source image $I^{S}$ and a target prompt $P$, PnP first performs DDIM inversion to obtain the image’s corresponding noise $\{\mathbf{z}^{S}_{t}\}_{t=1}^{T}$ at each time step $t$.
+It then collects the convolution features $\{f^{l}_{t}\}$ and attention scores $\{A^{l}_{t}\}$ from some predefined layers $l$ at each time step $t$ of the backward diffusion process $\mathbf{z}_{t-1}^{S}=\epsilon_{\theta}(\mathbf{z}_{t}^{S},\varnothing,t)$, where $\varnothing$ denotes the null text prompt during denoising.
+To generate the edited image $I^{*}$, PnP starts from the initial noise of the source image (i.e.
+$\mathbf{z}_{T}^{*}=\mathbf{z}_{T}^{S}$) and performs feature injection during denoising: $\mathbf{z}_{t-1}^{*}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},P,t,\{f^{l}_{t},A^{l}_{t}\})$, where $\epsilon_{\theta}(\cdot,\cdot,\cdot,\{f^{l}_{t},A^{l}_{t}\})$ represents the operation of replacing the intermediate feature and attention scores $\{f^{l*}_{t},A^{l*}_{t}\}$ with $\{f^{l}_{t},A^{l}_{t}\}$.
+This feature injection mechanism ensures $I^{*}$ to preserve the layout and structure from $I^{S}$ while reflecting the description in $P$.
+To control the feature injection strength, PnP also employs two thresholds $\tau_{f}$ and $\tau_{A}$ such that the feature and attention scores are only injected in the first $\tau_{f}$ and $\tau_{A}$ denoising steps.
+Our method extends this feature injection mechanism to I2V generation models, where we inject features in convolution, spatial, and temporal attention layers.
+We show the detailed design of AnyV2V in Section 4.
 4 AnyV2V
@@
 An overall illustration of our video editing pipeline is shown in Figure 2.
-We provide a discussion of the preliminaries of our method in Appendix C.
 In this section, we explain each core component of our method.
@@
 AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame.
-This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [15, 17, 34, 48], mask-based image editing [36, 1], image inpainting [45, 29], identity-preserving image editing [49], and subject-driven image editing [9, 30, 18].
+This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer (Gatys et al., 2015; Ghiasi et al., 2017; Lötzsch et al., 2022; Wang et al., 2024a), mask-based image editing (Nichol et al., 2022; Avrahami et al., 2022), image inpainting (Suvorov et al., 2021; Ku et al., 2022), identity-preserving image editing (Wang et al., 2024b), and subject-driven image editing (Chen et al., 2023c; Li et al., 2023; Gu et al., 2023a).
 4.2 Structural Guidance using DDIM Inverison
@@
 We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors.
-We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance [22].
+We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance (Ho & Salimans, 2022).
 We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
@@
 Therefore, it is likely that a large portion of the motion information is encoded in the temporal layers of the I2V generation models.
-Concurrent work [2] also observes that features in the temporal layers show similar characteristics with optical flow [26], a pattern that is often used to describe the motion of the video.
+Concurrent work (Bai et al., 2024) also observes that features in the temporal layers show similar characteristics with optical flow (Horn & Schunck, 1981), a pattern that is often used to describe the motion of the video.
 To better reconstruct the source video motion in the edited video, we propose to also inject the temporal attention features in the video generation process.
@@
 We employ AnyV2V on three off-the-shelf I2V generation models: I2VGen-XL [^1] We use the version provided in https://huggingface.co/ali-vilab/i2vgen-xl.
-[59], ConsistI2V [42] and SEINE [10].
+(Zhang et al., 2023c), ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d).
 For all I2V models, we use $\tau_{conv}=0.2T$, $\tau_{sa}=0.2T$ and $\tau_{ta}=0.5T$, where $T$ is the total number of sampling steps.
-We use the DDIM [44] sampler and set $T$ to the default values of the selected I2V models.
-Following PnP [47], we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections.
-During sampling, we apply text classifier-free guidance (CFG) [22] for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits.
-To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix [4], style transfer model Neural Style Transfer (NST) [15], subject-driven image editing model AnyDoor [9], and identity-driven image editing model InstantID [49].
+We use the DDIM (Song et al., 2020) sampler and set $T$ to the default values of the selected I2V models.
+Following PnP (Tumanyan et al., 2023b), we set $l_{1}=4$ for convolution feature injection and $l_{2}=l_{3}=\{4,5,6,...,11\}$ for spatial and temporal attention injections.
+During sampling, we apply text classifier-free guidance (CFG) (Ho & Salimans, 2022) for all models with the same negative prompt “Distorted, discontinuous, Ugly, blurry, low resolution, motionless, static, disfigured, disconnected limbs, Ugly faces, incomplete arms” across all edits.
+To obtain the initial edited frames in our implementation, we use a set of image editing model candidates including prompt-based image editing model InstructPix2Pix (Brooks et al., 2023), style transfer model Neural Style Transfer (NST) (Gatys et al., 2015), subject-driven image editing model AnyDoor (Chen et al., 2023c), and identity-driven image editing model InstantID (Wang et al., 2024b).
 We experiment with only the successfully edited frames, which is crucial for our method.
@@
 Alignment Overall CLIP-Text CLIP-Image
-Tune-A-Video [53] 15.2% 2.1% 0.2902 0.9704
+Tune-A-Video 15.2% 2.1% 0.2902 0.9704
 […]
@@
 AnyV2V (I2VGen-XL) 69.7% 46.2% 0.2932 0.9652
+5.2 Tasks Definition
+  • 1.
+    Prompt-based Editing: allows users to manipulate video content using only natural language.
+This can include descriptive prompts or instructions.
+With the prompt, Users can perform a wide range of edits, such as incorporating accessories, spawning or swapping objects, adding effects, or altering the background.
+  • 2.
+    Reference-based Style Transfer: In the realm of style transfer tasks, the artistic styles of Monet and Van Gogh are frequently explored, but in real-life examples, users might want to use a distinct style based on one particular artwork.
+In reference-based style transfer, we focus on using a style image as a reference to perform video editing.
+The edited video should capture the distinct style of the referenced artwork.
+  • 3.
+    Subject-driven Editing: In subject-driven video editing, we aim at replacing an object in the video with a target subject based on a given subject image while maintaining the video motion and persevering the background.
+  • 4.
+    Identity Manipulation: Identity manipulation allows the user to manipulate video content by replacing a person with another person’s identity in the video based on an input image of the target person.
 5.3 Evaluation Results
 As shown in Figure 3 and 4, AnyV2V can perform the following video editing tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation.
-we compare AnyV2V against three baseline models Tune-A-Video [53], TokenFlow [16] and FLATTEN [12] for (1) prompt-based editing.
+we compare AnyV2V against three baseline models Tune-A-Video (Wu et al., 2023b), TokenFlow (Geyer et al., 2023) and FLATTEN (Cong et al., 2023) for (1) prompt-based editing.
 Since there exists no publicly available baseline method for task (2) (3) (4), we evaluate the performance of three I2V generation models under AnyV2V.
 We included a more comprehensive evaluation in Appendix B.
-Prompt-based Editing Unlike the baseline methods [53, 16, 12] that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged.
-Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix [4], AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue.
+Prompt-based Editing
+Unlike the baseline methods (Wu et al., 2023b; Geyer et al., 2023; Cong et al., 2023) that often introduce unwarranted changes not specified in the text commands, AnyV2V utilizes the precision of image editing models to ensure only the targeted areas of the scene are altered, leaving the rest unchanged.
+Combining AnyV2V with the instruction-guided image editing model InstructPix2Pix (Brooks et al., 2023), AnyV2V accurately places a party hat on an elderly man’s head and correctly paints the plane in blue.
 Additionally, it maintains the original video’s background and fidelity, whereas, in comparison, baseline methods often alter the color tone and shape of objects, as illustrated in Figure 3.
@@
 For quantitative evaluations, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting, and also compute the metrics CLIP-Text for text alignment and CLIP-Image for temporal consistency.
-In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model [41] and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames.
+In detail, CLIP-Text is computed by the average cosine similarity between text embeddings from CLIP model (Radford et al., 2021) and CLIP-Image is computed in the same way but between image embeddings for every pair of consecutive frames.
 Table 2 shows that AnyV2V generally achieves high text-alignment and temporal consistency, while AnyV2V with I2VGen-XL backbone is the most preferred method because it does not edit the video precisely.
-Style Transfer, Subject-Driven Editing and Identity Manipulation For these novel tasks, we stress the alignment with the reference image instead of the text prompt.
+Style Transfer, Subject-Driven Editing and Identity Manipulation
+For these novel tasks, we stress the alignment with the reference image instead of the text prompt.
 As shown in Figure 4, we can observe that for task (2), AnyV2V can capture one particular style that is tailor-made, and even if the style is not learned by the text encoder.
@@
 We report both human evaluation results and find that AnyV2V with I2VGen-XL backbone is the most preferred method in terms of reference alignment and overall performance, plotted in Table 5, which is in Appendix B.
-I2V Backbones We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively.
+I2V Backbones
+We find that AnyV2V (I2VGen-XL) tends to be the most robust both qualitatively and quantitatively.
 It has a good generalization ability to produce consistent motions in the video with high visual quality.
@@
 AnyV2V (SEINE)’s generalization ability is relatively weaker but still produces consistent and high-quality video if the motion is simple enough, such as a person walking.
-Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models [10, 42, 59] are mostly trained on video data that contains only 16 frames.
-To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V [42] and SEINE [10].
-However, we find that such an experiment setup cannot maintain semantic consistency in our case.
-As many works in extending video generation exploit the initial latent to generate longer video [40, 54], we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output.
-Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
 Figure 5: AnyV2V can edit video length beyond the training frame while maintaining motion consistency.
 The first row is the source video frames while the second rows are the edited.
-The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix [4].
+The editing prompt of the image was “turn woman into a robot” using image model InstructPix2Pix (Brooks et al., 2023).
+Editing Video beyond Training Frames of I2V model Current state-of-the-art I2V models (Chen et al., 2023d; Ren et al., 2024; Zhang et al., 2023c) are mostly trained on video data that contains only 16 frames.
+To edit videos that have length beyond the training frames of the I2V model, an intuitive approach would be generating videos in an auto-regressive manner as used in ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d).
+However, we find that such an experiment setup cannot maintain semantic consistency in our case.
+As many works in extending video generation exploit the initial latent to generate longer video (Qiu et al., 2023; Wu et al., 2023c), we leverage the longer inverted latent as the initial latent and force an I2V model to generate longer frames of output.
+Our experiments found that the inverted latent contains enough temporal and semantic information to allow the generated video to maintain temporal and semantic consistency, as shown in Figure 5.
 5.4 Ablation Study
@@
 We feed in the DDIM inverted noise to the I2V models such that the generated videos (first row) are reconstructions of the source video.
-To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL [59], ConsistI2V [42] and SEINE [10].
+To better understand how different layers in the I2V denoising U-Net produce features during video sampling, we perform a visualization of the convolution, spatial and temporal attention features for the three candidate I2V models I2VGen-XL (Zhang et al., 2023c), ConsistI2V (Ren et al., 2024) and SEINE (Chen et al., 2023d).
 Specifically, we visualize the average activation values across all channels in the output feature map from the convolution layers, and the average attention scores across all attention heads and all tokens (i.e.
@@
 The experiment was conducted with the SEINE backbone.
-Appendix B Evaluation
-B.1 Tasks Definition
-  • 1.
-    Prompt-based Editing: allows users to manipulate video content using only natural language.
-This can include descriptive prompts or instructions.
-With the prompt, Users can perform a wide range of edits, such as incorporating accessories, spawning or swapping objects, adding effects, or altering the background.
-  • 2.
-    Reference-based Style Transfer: In the realm of style transfer tasks, the artistic styles of Monet and Van Gogh are frequently explored, but in real-life examples, users might want to use a distinct style based on one particular artwork.
-In reference-based style transfer, we focus on using a style image as a reference to perform video editing.
-The edited video should capture the distinct style of the referenced artwork.
-  • 3.
-    Subject-driven Editing: In subject-driven video editing, we aim at replacing an object in the video with a target subject based on a given subject image while maintaining the video motion and persevering the background.
-  • 4.
-    Identity Manipulation: Identity manipulation allows the user to manipulate video content by replacing a person with another person’s identity in the video based on an input image of the target person.
-B.2 Quantitative Evaluations
+Appendix B Evaluation Detail
+B.1 Quantitative Evaluations
 Prompt-based Editing
 For (1) prompt-based editing, we conduct a human evaluation to examine the degree of prompt alignment and overall preference of the edited videos based on user voting.
-We compare AnyV2V against three baseline models: Tune-A-Video [53], TokenFlow [16] and FLATTEN [12].
+We compare AnyV2V against three baseline models: Tune-A-Video (Wu et al., 2023b), TokenFlow (Geyer et al., 2023) and FLATTEN (Cong et al., 2023).
 Human evaluation results in Table 2 demonstrate that our model achieves the best overall preference and prompt alignment among all methods, and AnyV2V (I2VGen-XL) is the most preferred method.
@@
 We also employ automatic evaluation metrics on our edited video of the human evaluation datasets.
-Following previous works [5, 2], our automatic evaluation employs the CLIP [41] model to assess both text alignment and temporal consistency.
+Following previous works (Ceylan et al., 2023; Bai et al., 2024), our automatic evaluation employs the CLIP (Radford et al., 2021) model to assess both text alignment and temporal consistency.
 For text alignment, we calculate the CLIP-Score, specifically by determining the average cosine similarity between the CLIP text embeddings derived from the editing prompt and the CLIP image embeddings across all frames.
@@
 On the other hand, AnyV2V (ConsistI2V) shines in subject-driven editing, achieving second-best results in both reference alignment and overall preference.
-Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing [28], we also report the image editing success rate in Table 5 to clarify that our method relies on a good image frame edit.
-B.3 Qualitative Results
+Since the latest image editing models have not yet reached a level of maturity that allows for consistent and precise editing (Ku et al., 2024), we also report the image editing success rate in Table 5 to clarify that our method relies on a good image frame edit.
+B.2 Qualitative Results
 Prompt-based Editing
 By leveraging the strength of image editing models, our AnyV2V framework provides precise control of the edits such that the irrelevant parts in the scene are untouched after editing.
-In our experiment, we used InstructPix2Pix [4] for the first frame edit.
+In our experiment, we used InstructPix2Pix (Brooks et al., 2023) for the first frame edit.
 Shown in Figure 3, our method correctly places a party hat on an old man’s head and successfully turns the color of an airplane to blue, while preserving the background and keeping the fidelity to the source video.
-Comparing our work with the three baseline models TokenFlow [16], FLATTEN [12], and Tune-A-Video [53], the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt.
+Comparing our work with the three baseline models TokenFlow (Geyer et al., 2023), FLATTEN (Cong et al., 2023), and Tune-A-Video (Wu et al., 2023b), the baseline methods display either excessive or insufficient changes in the edited video to align with the editing text prompt.
 The color tone and object shapes are also tilted.
@@
 Reference-based Style Transfer
-Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST [15] to obtain the edited frame.
+Our approach diverges from relying solely on textual descriptors for conducting style edits, using the style transfer model NST (Gatys et al., 2015) to obtain the edited frame.
 This level of controllability offers artists the unprecedented opportunity to use their art as a reference for video editing, opening new avenues for creative expression.
@@
 Subject-driven Editing
-In our experiment, we employed a subject-driven image editing model AnyDoor [9] for the first frame editing.
+In our experiment, we employed a subject-driven image editing model AnyDoor (Chen et al., 2023c) for the first frame editing.
 AnyDoor allows replacing any object in the target image with the subject from only one reference image.
@@
 Identity Manipulation
-By integrating the identity-preserved image personalization model InstantID [49] with ControlNet [58], this approach enables the replacement of an individual’s identity to create an initial frame.
+By integrating the identity-preserved image personalization model InstantID (Wang et al., 2024b) with ControlNet (Zhang et al., 2023b), this approach enables the replacement of an individual’s identity to create an initial frame.
 Our AnyV2V framework then processes this initial frame to produce an edited video, swapping the person’s identity as showcased in Figure 4.
@@
 It is possible to leverage other identity-preserved image personalization models and apply them to AnyV2V to preserve the background.
-B.4 Human Evaluation
-B.4.1 Dataset
+B.3 Human Evaluation
+B.3.1 Dataset
 Our human evaluation dataset contains a total of 89 samples that have been collected from https://www.pexels.com.
-For prompt-based editing, we employed InstructPix2Pix [4] to compose the examples.
+For prompt-based editing, we employed InstructPix2Pix (Brooks et al., 2023) to compose the examples.
 Topics include swapping objects, adding objects, and removing objects.
-For subject-driven editing, we employed AnyDoor [9] to replace objects with reference subjects.
-For Neural Style Transfer, we employed NST [15] to compose the examples.
-For identity manipulation, we employed InstantID [49] to compose the examples.
+For subject-driven editing, we employed AnyDoor (Chen et al., 2023c) to replace objects with reference subjects.
+For Neural Style Transfer, we employed NST (Gatys et al., 2015) to compose the examples.
+For identity manipulation, we employed InstantID (Wang et al., 2024b) to compose the examples.
 See Table 4 for the statistic.
@@
 Category Number of Entries Image Editing Models Used
-Prompt-based Editing 45 InstructPix2Pix [4]
-Reference-based Style Transfer 20 NST [15]
-Identity Manipulation 13 InstantID [49]
-Subject-driven Editing 11 AnyDoor [9]
+Prompt-based Editing 45 InstructPix2Pix
+Reference-based Style Transfer 20 NST
+Identity Manipulation 13 InstantID
+Subject-driven Editing 11 AnyDoor
 Total 89
-B.4.2 Interface
+B.3.2 Interface
 In the evaluation setup, we provide the evaluator with generated images from both the baseline models and the AnyV2V models.
@@
 Figure 10: The interface of individual evaluation.
-Appendix C Preliminary
-C.1 Image-to-Video (I2V) Generation Models
-In this work, we focus on leveraging latent diffusion-based [43] I2V generation models for video editing.
-Given an input first frame $I_{1}$, a text prompt $\mathbf{s}$ and a noisy video latent $\mathbf{z}_{t}$ at time step $t$, I2V generation models recover a less noisy latent $\mathbf{z}_{t-1}$ using a denoising model $\epsilon_{\theta}(\mathbf{z}_{t},I_{1},\mathbf{s},t)$ conditioned on both $I_{1}$ and $\mathbf{s}$.
-The denoising model $\epsilon_{\theta}$ contains a set of spatial and temporal self-attention layers, where the self-attention operation can be formulated as:
-$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (3)
-$\mathrm{Attention}(Q,K,V)=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})V,$ (4)
-where $z$ is the input hidden state to the self-attention layer and $W^{Q}$, $W^{K}$ and $W^{V}$ are learnable projection matrices that map $z$ onto query, key and value vectors, respectively.
-For spatial self-attention, $z$ represents a sequence of spatial tokens from each frame.
-For temporal self-attention, $z$ is composed of tokens located at the same spatial position across all frames.
-C.2 DDIM Inversion
-The denoising process for I2V generation models from $\mathbf{z}_{t}$ to $\mathbf{z}_{t-1}$ can be achieved using the DDIM [44] sampling algorithm.
-The reverse process of DDIM sampling, known as DDIM inversion [35, 13], allows obtaining $\mathbf{z}_{t+1}$ from $\mathbf{z}_{t}$ such that $\mathbf{z}_{t+1}=\sqrt{\frac{\alpha_{t+1}}{\alpha_{t}}}\mathbf{z}_{t}+(\sqrt{\frac{1}{\alpha_{t+1}}-1}-\sqrt{\frac{1}{\alpha_{t}}-1})\cdot\epsilon_{\theta}(\mathbf{z}_{t},x_{0},\mathbf{s},t)$, where $\alpha_{t}$ is derived from the variance schedule of the diffusion process.
-C.3 Plug-and-Play (PnP) Diffusion Features
-Tumanyan et al.
-[46] proposed PnP diffusion features for image editing, based on the observation that intermediate convolution features $f$ and self-attention scores $A=\mathrm{Softmax}(\frac{QK^{\top}}{\sqrt{d}})$ in a text-to-image (T2I) denoising U-Net capture the semantic regions (e.g.
-legs or torso of a human body) during the image generation process.
-Given an input source image $I^{S}$ and a target prompt $P$, PnP first performs DDIM inversion to obtain the image’s corresponding noise $\{\mathbf{z}^{S}_{t}\}_{t=1}^{T}$ at each time step $t$.
-It then collects the convolution features $\{f^{l}_{t}\}$ and attention scores $\{A^{l}_{t}\}$ from some predefined layers $l$ at each time step $t$ of the backward diffusion process $\mathbf{z}_{t-1}^{S}=\epsilon_{\theta}(\mathbf{z}_{t}^{S},\varnothing,t)$, where $\varnothing$ denotes the null text prompt during denoising.
-To generate the edited image $I^{*}$, PnP starts from the initial noise of the source image (i.e.
-$\mathbf{z}_{T}^{*}=\mathbf{z}_{T}^{S}$) and performs feature injection during denoising: $\mathbf{z}_{t-1}^{*}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},P,t,\{f^{l}_{t},A^{l}_{t}\})$, where $\epsilon_{\theta}(\cdot,\cdot,\cdot,\{f^{l}_{t},A^{l}_{t}\})$ represents the operation of replacing the intermediate feature and attention scores $\{f^{l*}_{t},A^{l*}_{t}\}$ with $\{f^{l}_{t},A^{l}_{t}\}$.
-This feature injection mechanism ensures $I^{*}$ to preserve the layout and structure from $I^{S}$ while reflecting the description in $P$.
-To control the feature injection strength, PnP also employs two thresholds $\tau_{f}$ and $\tau_{A}$ such that the feature and attention scores are only injected in the first $\tau_{f}$ and $\tau_{A}$ denoising steps.
-Our method extends this feature injection mechanism to I2V generation models, where we inject features in convolution, spatial, and temporal attention layers.
-We show the detailed design of AnyV2V in Section 3.
 […]
@@
 Style Transfer Editing Manipulation
-Image Editing Method NST [15] AnyDoor [9] InstantID [49]
+Image Editing Method NST AnyDoor InstantID
 Image Editing Success Rate $\approx$90$\approx$10$\approx$80%
@@
 AnyV2V (I2VGen-XL) 100.0% 76.9% 93.9% 84.8% 90.1% 45.4%
-Appendix D Discussion
-D.1 Limitations
+Appendix C Discussion
+C.1 Limitations
 Inaccurate Edit from Image Editing Models.
 As our method relies on an initial frame edit, the image editing models are used.
-However, the current state-of-the-art models are not mature enough to perform accurate edits consistently [28].
-For example, in the subject-driven video editing task, we found that AnyDoor [9] requires several tries to get a good editing result.
+However, the current state-of-the-art models are not mature enough to perform accurate edits consistently (Ku et al., 2024).
+For example, in the subject-driven video editing task, we found that AnyDoor (Chen et al., 2023c) requires several tries to get a good editing result.
 Efforts are required in manually picking a good edited frame.
@@
 We anticipate that the presence of a robust I2V model can address this issue.
-D.2 License of Assets
-For image editing models, InstructPix2Pix [4] inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion.
-Neural Style Transfer [15] is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License.
-InstantID [49] is under Apache License 2.0.
-AnyDoor [9] is under the MIT License.
-For baselines, Tune-A-Video [53] is under Apache License 2.0, TokenFlow [16] is under MIT License andFLATTEN [12] is under Apache License 2.0.
+C.2 License of Assets
+For image editing models, InstructPix2Pix (Brooks et al., 2023) inherits Creative ML OpenRAIL-M License as it is built upon Stable Diffusion.
+Neural Style Transfer (Gatys et al., 2015) is under Creative Commons Attribution 4.0 License, and code samples are licensed under the Apache 2.0 License.
+InstantID (Wang et al., 2024b) is under Apache License 2.0.
+AnyDoor (Chen et al., 2023c) is under the MIT License.
+For baselines, Tune-A-Video (Wu et al., 2023b) is under Apache License 2.0, TokenFlow (Geyer et al., 2023) is under MIT License andFLATTEN (Cong et al., 2023) is under Apache License 2.0.
 For the human evaluation dataset, the dataset has been collected from https://www.pexels.com, with all data governed by the terms outlined at https://www.pexels.com/license/.
 We decide to release AnyV2V code under the Creative Commons Attribution 4.0 License for easy access in the research community.
-D.3 Societal Impacts
+C.3 Societal Impacts
 Postive Social Impact.
@@
 To mitigate the risk of misuse, one proposed solution is the adoption of unseen watermarking, a method commonly used to tackle such concerns in image generation.
-D.4 Safeguards
+C.4 Safeguards
 It is crucial to implement proper safeguards and responsible AI frameworks when developing user-friendly video editing systems.
@@
 We only collect images that are considered safe.
-D.5 Ethical Concerns for Human Evaluation
+C.5 Ethical Concerns for Human Evaluation
 We believe our proposed human evaluation does not incur ethical concerns due to the following reasons: (1) the study does not involve any form of intervention or interaction that could affect the participants’ well-being.
@@
 (4) Participation in the study is entirely voluntary, and participants can withdraw at any time without any consequences.
-D.6 New Assets
+C.6 New Assets
 Our paper introduces several new assets including a human evaluation dataset and demo videos generated by AnyV2V.
@@
 All assets are shared under an open license to facilitate reuse and further research.
-Appendix E More AnyV2V Showcases
-Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix [4].
+Appendix D More AnyV2V Showcases
+Figure 11: AnyV2V becomes an instruction-based video editing tool when plugged with instruction-guided image editing models like InstructPix2Pix (Brooks et al., 2023).
 Prompt used “Turn the couple into robots”, “Turn horse into zebra”, and “Turn the sand into snow”.
-Figure 12: The more recent model InstantStyle [48] can seamlessly plug in AnyV2V to perform reference-based style transfer video editing.
+Figure 12: The more recent model InstantStyle (Wang et al., 2024a) can seamlessly plug in AnyV2V to perform reference-based style transfer video editing.
 The reference art image in the bottom left corner is used to retrieve the first image edit.
-Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor [9].
+Figure 13: AnyV2V can perform subject-driven video editing with a single image reference, using a subject-driven image editing model like AnyDoor (Chen et al., 2023c).
 The subject image in the bottom left corner is used to retrieve the first image edit.